### v2

# A globally convergent gradient method with momentum

Abstract
In this work, we consider smooth unconstrained optimization problems and we deal with the class of gradient methods with momentum, i.e., descent algorithms where the search direction is defined as a linear combination of the current gradient and the preceding search direction. This family of algorithms includes nonlinear conjugate gradient methods and Polyak’s heavy-ball approach, and is thus of high practical and theoretical interest in large-scale nonlinear optimization. We propose a general framework where the scalars of the linear combination defining the search direction are computed simultaneously by minimizing the approximate quadratic model in the 2 dimensional subspace. This strategy allows us to define a class of gradient methods with momentum enjoying global convergence guarantees and an optimal worst-case complexity bound in the nonconvex setting. Differently than all related works in the literature, the convergence conditions are stated in terms of the Hessian matrix of the bi-dimensional quadratic model. To the best of our knowledge, these results are novel to the literature. Moreover, extensive computational experiments show that the gradient methods with momentum here presented outperform classical conjugate gradient methods and are (at least) competitive with the state-of-art method for unconstrained optimization, i.e, L-BFGS method.
keywords:
Nonconvex optimization, momentum, global convergence, complexity bound
pacs:
MSC Classification]90C26, 90C30, 90C60
1 Introduction
In this work we consider unconstrained optimization problems
$\min_{x\in\mathbb{R}^{n}}f(x),$
where $f:\mathbb{R}^{n}\to\mathbb{R}$ is a continuously differentiable objective function. We do not assume that the function is convex. We focus on first order descent methods that exploit information from the preceding iteration to determine the search direction and the stepsize at the current one. We will hence refer to gradient methods with momentum, i.e., to algorithms defined by an iteration of the generic form
$x_{k+1}=x_{k}-\alpha_{k}\nabla f(x_{k})+\beta_{k}(x_{k}-x_{k-1}),$ (1)
where $\alpha_{k}>0$ is the stepsize, and $\beta_{k}>0$ is the momentum weight. Partially repeating the previous step has the effect of controlling oscillation and providing acceleration in low curvature regions. All of this can, in principle, be achieved only exploiting already available information: no additional function evaluations are required to be carried out. This feature makes the addition of momentum terms appealing in large-scale settings and, in particular, in the deep learning context [1, 2].
The best-known and most important gradient methods with momentum arguably are:
    Polyak’s heavy-ball method [3, 4];
    conjugate gradient methods (see, e.g., [5]).
Conjugate gradient methods can be described by the updates
$x_{k+1}=x_{k}+\alpha_{k}d_{k},\qquad d_{k+1}=-\nabla f(x_{k+1})+\beta_{k+1}d_{k},$
where $\alpha_{k}$ is computed by means of a line search, whereas $\beta_{k}$ is obtained according to one of many rules from the literature (see, e.g., [5]). The update of conjugate gradient methods can be rewritten as
$x_{k+1}$ $=x_{k}+\alpha_{k}(-\nabla f(x_{k})+\beta_{k}d_{k-1})$
$=x_{k}+\alpha_{k}\left(-\nabla f(x_{k})+\frac{\beta_{k}}{\alpha_{k-1}}(x_{k}-x_{k-1})\right)$
$=x_{k}-\alpha_{k}\nabla f(x_{k})+\hat{\beta}_{k}(x_{k}-x_{k-1}),$
so that it can be viewed as a gradient method with momentum according to definition (1).
The convergence theory of nonlinear conjugate gradient methods has been a research topic for about 30 years and now it can be considered well-established, while some complexity results have been stated only recently [6, 7]. Several proposed conjugate gradient methods can be considered sound and efficient tools for unconstrained optimization. Recently, a new class of conjugate gradient methods, known as subspace minimization conjugate gradient (SMCG) methods, have been proposed in the literature [8, 9, 10, 11]. We will discuss more in detail this class of approaches later in this work, since they are related to the framework proposed here.
On the other hand, the heavy-ball algorithm is directly described by an iteration of the form (1), where $\alpha_{k}$ and $\beta_{k}$ typically are fixed positive values[12]; in principle, suitable constants should be chosen depending on properties of the objective function (e.g., using Lipschitz constant of the gradient or the constant of strong convexity) [13, 14, 4]. In practice, however, this information is often not accessible and thus $\alpha_{k}$ can be chosen by a line search while the momentum parameter $\beta_{k}$ is usually blindly set to some (more or less) reasonable value. Convergence results for the heavy-ball method have been proven in the convex case [13, 3, 15], while the convergence of the method in the nonconvex case is still an open problem.
Thus, algorithmic issues related to the choice of the two parameters $\alpha_{k}$ and $\beta_{k}$ and the theoretical gap related to the convergence in the nonconvex case did not allow, until now, to include the heavy-ball method within the class of sound methods for smooth unconstrained optimization. As a matter of fact, there does not exist any popular software implementation of the method. However, heavy-ball type momentum terms are consistently and effectively used within modern frameworks of stochastic optimization for neural network training [16, 17].
Both conjugate gradient and heavy-ball methods are therefore of practical interest in large-scale nonlinear optimization settings. We hence believe it is worth to focus on the study of the general class of gradient methods with momentum, in order to possibly define convergent algorithms improving the efficiency of standard nonlinear conjugate gradient methods.
To this aim, we draw inspiration from the idea presented in [11], where the search direction is computed by minimizing the approximate quadratic model in the 2 dimensional subspace spanned by the current gradient and the last search direction. According to this approach, the scalars $\alpha_{k}$ and $\beta_{k}$ are not prefixed, but rather they are simultaneously determined by a bidimensional search. We define a general framework of gradient methods with momentum based on a simple Armijo-type line search, and we prove global convergence results under the first-order smoothness assumptions only. We also derive specific algorithms with momentum from the general framework. Furthermore, we provide complexity results, proving for the proposed gradient method with momentum the worst-case complexity bound of ${\cal O}(\epsilon^{-2})$, which is optimal for first order algorithms in the nonconvex setting [18]. Up to our knowledge, the presented algorithm is the first framework of gradient methods with momentum having in the nonconvex case both theoretical convergence properties and a complexity bound.
Extensive computational experiments show that the gradient methods with momentum here presented outperform conjugate gradient methods and are (at least) competitive with the state-of-art method for unconstrained optimization, i.e, L-BFGS method.
The rest of the paper is organized as follows: we describe the main idea of the work in section 2, with a focus on some important related works in section 2.1; in section 3 we discuss conditions for the proposed method to be well defined. Then, we discuss in section 4 the issue of guaranteeing that the employed search directions are gradient related; in particular, we consider the cases where the property descends from assumptions on either $n\times n$ or $2\times 2$ matrices (section 4.1 and section 4.2 respectively). In section 5, we finally describe the proposed algorithmic framework for gradient methods with momentum, formalizing convergence and complexity results. In section 6, we then discuss concrete strategies to estimate the matrices that are at the core of the proposed method. In section 7, we report the results of thorough computational experiments empirically showing the potential of the proposed class of approaches. We finally give some concluding remarks in section 8.
2 The main idea
A vast class of iterative algorithms for nonlinear unconstrained optimization (namely linesearch algorithms) can be written in a general form as
$x_{k+1}=x_{k}+\alpha_{k}d_{k},$
where $d_{k}\in\mathbb{R}^{n}$ is the search direction and $\alpha_{k}>0$ is the stepsize. The most typical rules for the choice of the direction follow a general scheme, which is basically given by the following optimization subproblem:
$\min_{d\in\mathbb{R}^{n}}\;\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}B_{k}d,$ (2)
where $B_{k}$ is a suitable symmetric positive definite matrix. By properly choosing $B_{k}$, we retrieve standard methods; in particular:
  • if $B_{k}=I$, then we obtain the steepest descent direction $-\nabla f(x_{k})$;
  • if $B_{k}=\nabla^{2}f(x_{k})$, then we obtain Newton’s direction $-\left[\nabla^{2}f(x_{k})\right]^{-1}\nabla f(x_{k})$;
  • if $B_{k}$ is a positive definite matrix obtained with suitable update rules, we obtain standard Quasi-Newton updates.
Gradient methods with momentum can be considered in the above framework by adding into (2) a suitable constraint on $d$, i.e.,
$\min_{d,\alpha,\beta}$ $g_{k}^{T}d+\frac{1}{2}d^{T}B_{k}d$ (3)
s.t. $d=-\alpha g_{k}+\beta s_{k},$
where $g_{k}=\nabla f(x_{k})$ and $s_{k}=x_{k}-x_{k-1}$. Then, substituting the constraint into the objective function of (3), the problem reduces to
$\min_{\alpha,\beta}\phi(\alpha,\beta)$ (4)
where
$\phi(\alpha,\beta)=-\alpha\|g_{k}\|^{2}+\beta g_{k}^{T}s_{k}+\frac{1}{2}\alpha^{2}g_{k}^{T}B_{k}g_{k}+\frac{1}{2}\beta^{2}s_{k}^{T}B_{k}s_{k}-\alpha\beta g_{k}^{T}B_{k}s_{k},$
or equivalently
$\phi(\alpha,\beta)=\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (5)
We can denote the $2\times 2$ matrix in the above equation as
$H_{k}=\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}=\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix}=P_{k}^{T}B_{k}P_{k},$ (6)
where we omit the dependence of $H_{ij}$ on $k$ for the sake of notation simplicity and we set $P_{k}=\left[-g_{k}\;\;s_{k}\right]$. Letting $u=[\alpha\;\;\beta]^{T}$, we can express the problem as
$\min_{u}\;\frac{1}{2}u^{T}P_{k}^{T}B_{k}P_{k}u+g_{k}^{T}P_{k}u,$ (7)
or equivalently as
$\min_{u}\;\frac{1}{2}u^{T}H_{k}u+g_{k}^{T}P_{k}u.$ (8)
Once a solution $u_{k}=[\alpha_{k}\;\;\beta_{k}]^{T}$ of (7) is determined, we define
$d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$
and, provided that $d_{k}$ is a descent direction, we set
$x_{k+1}=x_{k}+\eta_{k}d_{k},$ (9)
where $\eta_{k}$ can be determined, for example, by an Armjio-type line search.
Now consider the two-dimensional function
$\psi_{k}(\alpha,\beta)=f(x_{k}-\alpha g_{k}+\beta s_{k})$
and assume that $f$ is twice continuously differentiable. We have
$\psi_{k}(0,0)=f(x_{k}),\qquad\nabla\psi_{k}(0,0)=\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix},$
$\nabla^{2}\psi_{k}(0,0)=\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}.$
The quadratic Taylor polynomial for $\psi_{k}(\alpha,\beta)$ centered at $(0,0)$ is thus
$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (10)
The matrix $H_{k}$ defined by (6) can then be viewed as a matrix approximating $\nabla^{2}\psi_{k}(0,0)$. From this point of view, $H_{k}$ could in fact be any $2\times 2$ matrix (independent of the $n\times n$ matrix $B_{k}$) related to the approximation of the quadratic Taylor polynomial of $\psi_{k}(\alpha,\beta)=f(x_{k}-\alpha g_{k}+\beta s_{k})$, i.e.,
$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (11)
Setting $u=[\alpha\;\;\beta]^{T}$, the minimization problem of (11) leads again to problem (8), being $H_{k}$ a generic $2\times 2$ matrix with suitable properties.
From a theoretical point of view, the following issues must me considered:
  1. (a)
    we shall state conditions ensuring that the two-dimensional subproblem (7) admits solution;
  2. (b)
    we shall analyze under which conditions the obtained search direction $d_{k}$, coupled with line search techniques, allows to prove global convergence properties and complexity results for the iterative scheme (9).
In the following, we will state conditions concerning either the $n\times n$ matrix $B_{k}$ appearing in (7) or directly the $2\times 2$ matrix $H_{k}$ in (8) (with no explicit connection to $B_{k}$), in order to satisfy the above theoretical requirements.
2.1 Related Works
We now briefly analyze some important works directly related to our approach.
Subspace minimization CG
Subspace minimization conjugate gradient (SMCG) methods, for which we refer the reader, for example, to the very recent paper [8], consider the two-dimensional subproblem (4)-(5) by assuming that the $n\times n$ matrix $B_{k}$ satisfies the secant equation
$B_{k}s_{k}=g_{k}-g_{k-1}.$
The issue deserving attention becomes that of suitably managing the term $g_{k}^{T}B_{k}g_{k}=\rho_{k}$. Several strategies have been proposed in connection with this issue. Some global convergence results are stated, while complexity results are not known. However, as clearly written in [8], “A question is naturally to be asked: can one develop an efficient SMCG method without determining the parameter $\rho_{k}$?”. By our work, we give a positive answer to this question.
Common-directions methods
The framework discussed in [19] considers (7)-(9) in a more general setting, i.e., by assuming that $P_{k}$ is a general $n\times m_{k}$ matrix, with $m_{k}\geq 1$, containing at least a search direction which satisfies an angle condition. The columns of $P_{k}$ are required to be linearly independent. Convergence results are stated by assuming that $\{B_{k}\}$ is a sequence of uniformly positive definite symmetric matrices bounded above. Then, gradient methods with momentum fits this framework provided that $g_{k}$ and $s_{k}$ are linearly independent. The convergence analysis relies on the properties of the $n\times n$ matrix $B_{k}$.
(Riemannian) dimension-reduced second-order method (RDRSOM)
The basic idea of this method, proposed in [20] and then extended to optimization on Riemannian manifolds [21], is similar to that of [19]. This time, the search direction is determined by solving a cubic-regularized subproblem on the selected subspace which, in the simplest case, is the two-dimensional subspace spanned by the gradient and the previous direction. The iteration complexity of the proposed algorithm is ${\cal O}(\epsilon^{-3/2})$. However, as remarked by the same authors, this complexity result is based on a very restrictive assumption requiring that the Hessian matrix is approximated accurately enough in the considered subspace, which therefore should be adaptively selected. In order to overcome this nontrivial issue, a practical version of the algorithm is presented with iteration complexity ${\cal O}(\epsilon^{-2})$. Differently from [19], the convergence analysis relies on the properties the $2\times 2$ matrix $H_{k}$ and requires that the Hessian matrix of the objective function is Lipschitz continuous. This assumption can not be relaxed since the proposed method is based on the cubic regularization approach.
Our main theoretical contribution, with respect to all the works on SMCG methods and to [19], concerns the definition of conditions on the $2\times 2$ matrix $H_{k}$ sufficient to ensure that the sequence of search directions $\{d_{k}\}$ is gradient-related and hence to state convergence and complexity results for the proposed gradient methods with momentum. As regards the contribution with respect to [21], we do not state assumptions on the Hessian matrix of the objective function: we only require that the gradient is Lipschitz continuous. Thus, we attain the same complexity result under much weaker assumptions. We remark that the theoretical focus on the low-dimensional matrix $H_{k}$ is fundamental to design convergent algorithms for large-scale optimization regardless of the $n\times n$ matrix $B_{k}$. This issue yields relevant implications, deeply analyzed and discussed later, from theoretical, algorithmic and computational point of views.
3 Existence of a solution for the two dimensional problem
In this section we state conditions on the matrices $B_{k}$ and $H_{k}$ sufficient to ensure that the subproblem (8) admits solution. The condition on $H_{k}$ is rather simple to state.
Proposition 1.
Consider problem (8). If $H_{k}$ is a symmetric positive definite matrix, then problem (8) admits an optimal solution.
Proof.
The quadratic objective function is strictly convex and hence the problem admits a unique solution. ∎
Now, assume $H_{k}$ is obtained, starting from an $n\times n$ matrix $B_{k}$, according to (6). We can state the following condition.
Proposition 2.
Consider problem (3). If $g_{k}$ and $s_{k}$ are non-zero vectors and $B_{k}$ is a symmetric positive definite matrix, then problem (3) admits an optimal solution.
Proof.
Let $H_{k}$ be the matrix defined in (6). Let us consider problem (8) as equivalent reformulation of problem (3), i.e.,
$\min_{u}q(u)=$ $\frac{1}{2}u^{T}P_{k}^{T}B_{k}P_{k}u+g_{k}^{T}P_{k}u.$
Let us first suppose that $g_{k}$ and $s_{k}$ are linearly independent. By definition, $H_{k}=P_{k}^{T}B_{k}P_{k}$. Since $B_{k}$ is positive definite, and $P_{k}=[-g_{k}\quad s_{k}]$ is full rank, $H_{k}$ is positive definite; then, by proposition 1, problem (8) admits solution.
On the other hand, let us assume that $g_{k}$ and $s_{k}$ are linearly dependent. In this case, we can write $s_{k}=\sigma g_{k}$ and it results that the Hessian matrix of the quadratic function $q(u)$ is symmetric positive semi-definite. However, we can show that system
$P_{k}^{T}B_{k}P_{k}u=-P_{k}^{T}g_{k}$ (12)
admits solution, i.e., that there exists at least a point $\bar{u}$ such that $\nabla q(\bar{u})=0$, and hence that $\bar{u}$ is a global minimizer of the convex function $q(u)$. Indeed, we have
$P_{k}^{T}B_{k}P_{k}=\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-\sigma g_{k}^{T}B_{k}g_{k}\\ -\sigma g_{k}^{T}B_{k}g_{k}&\sigma^{2}g_{k}^{T}B_{k}g_{k}\end{bmatrix},\qquad-P_{k}^{T}g_{k}=\begin{bmatrix}\|g_{k}\|^{2}\\ -\sigma\|g_{k}\|^{2}\end{bmatrix}.$
By the positive definiteness of $B_{k}$, we have that $g_{k}^{\top}B_{k}g_{k}>0$ and we can then write that
$1=\text{rank}\left(P_{k}^{T}B_{k}P_{k}\right)=\text{rank}\left(\left[P_{k}^{T}B_{k}P_{k}\quad-P_{k}^{T}g_{k}\right]\right).$
This implies that system (12) admits a solution, which is also a solution of problem (8) and thus (3). ∎
4 Properties of the search directions
In this section we consider again subproblem (3) by means of the equivalent reformulations (7)-(8) and we assume that it admits solution $u_{k}=[\alpha_{k}\;\;\beta_{k}]^{T}$, i.e., that at least one assumption of the preceding section either on $B_{k}$ or on $H_{k}$ holds. We are interested in studying further conditions on the above matrices to ensure that the obtained search directions
$d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$
are gradient-related according to the following well-known definition [22, 23].
Definition 1.
A sequence of search directions $\{d_{k}\}$ is gradient-related to the sequence of solutions $\{x_{k}\}$ if there exist $c_{1}>0$ and $c_{2}>0$ such that, for all $k$, we have
$g_{k}^{T}d_{k}\leq-c_{1}\|g_{k}\|^{2},\qquad\|d_{k}\|\leq c_{2}\|g_{k}\|.$ (13)
As already said, the above property is a requirement sufficient to guarantee that the sequence generated according to the following scheme
$x_{k+1}=x_{k}+\eta_{k}d_{k},$
where $\eta_{k}$ is the stepsize computed by an Armjio-type line search, is globally convergent to stationary points, Moreover, under this assumption the algorithm can be proved to have a worst case iteration and function evaluations complexity bound of ${\cal O}(\epsilon^{-2})$ to reach a solution $\bar{x}$ with $\|\nabla f(\bar{x})\|\leq\epsilon$. Note that this bound is actually tight for first-order methods under standard first-order smoothness assumptions.
4.1 Conditions on ${B_{k}}$
First we state the following result concerning the $n\times n$ matrix $B_{k}$. Note that, from here onward, we will denote by $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ the minimum and maximum eigenvalues respectively of a symmetric matrix $A$.
Proposition 3.
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of (7).
Let $\{B_{k}\}\subseteq\mathbb{R}^{n\times n}$ be the sequence of symmetric matrices defining problems (7) and assume that there exist scalars $0<\eta_{1}\leq\eta_{2}$ such that
$\eta_{1}\leq\lambda_{\min}(B_{k})\leq\lambda_{\max}(B_{k})\leq\eta_{2}$ (14)
holds for all $k$. Then the direction $d_{k}$ satisfies the following conditions:
$g_{k}^{T}d_{k}\leq-\frac{\eta_{1}}{\eta_{2}^{2}}\|g_{k}\|^{2}$ (15)
$\frac{1}{\eta_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{\eta_{1}}\|g_{k}\|.$ (16)
Proof.
Let us recall problem (7)
$\min_{u}$ $\frac{1}{2}u^{T}P_{k}^{T}B_{k}P_{k}u+g_{k}^{T}P_{k}u,$
where $d=P_{k}u$, $P_{k}=[-g_{k}\quad s_{k}]$ and $u=[\alpha\;\beta]^{T}$.
From proposition 2 we have that there exists at least a solution $u_{k}$ of problem (3), satisfying the following linear system
$P_{k}^{T}B_{k}P_{k}u_{k}=-P_{k}^{T}g_{k}.$ (17)
Multiplying both the members in (17) by $u_{k}^{T}$, we obtain $u_{k}^{T}P_{k}^{T}B_{k}P_{k}u_{k}=-u_{k}^{T}P_{k}^{T}g_{k}$, i.e.,
$d_{k}^{T}B_{k}d_{k}=-g_{k}^{T}d_{k},$ (18)
being $d_{k}=P_{k}u_{k}$. Now, consider the first row of system (17), i.e.,
$-g_{k}^{T}B_{k}d_{k}=\|g_{k}\|^{2}.$
Using (18) and recalling assumption (14), we can write
$\|g_{k}\|^{2}=|g_{k}^{T}B_{k}d_{k}|\leq\|g_{k}\|\|B_{k}\|\|d_{k}\|\leq\lambda_{\max}(B_{k})\|g_{k}\|\|d_{k}\|\leq\eta_{2}\|g_{k}\|\|d_{k}\|,$
and hence we obtain
$\|d_{k}\|\geq{\frac{1}{\eta_{2}}}\|g_{k}\|.$ (19)
Recalling (18), using (14) and (19), it follows
$-g_{k}^{T}d_{k}\geq\lambda_{\min}(B_{k})\|d_{k}\|^{2}\geq\eta_{1}\|d_{k}\|^{2}\geq{\frac{\eta_{1}}{\eta_{2}^{2}}}\|g_{k}\|^{2}.$ (20)
Considering again (18) we can also write
$\|g_{k}\|\|d_{k}\|\geq-g_{k}^{T}d_{k}=d_{k}^{T}B_{k}d_{k}\geq\lambda_{\min}(B_{k})\|d_{k}\|^{2}\geq\eta_{1}\|d_{k}\|^{2},$
so that we have
$\|d_{k}\|\leq{\frac{1}{\eta_{1}}}\|g_{k}\|.$ (21)
Then, (19), (20) and (21) prove the thesis of the proposition. ∎
4.2 Conditions on $H_{k}$
In the previous section, problem (3) was rewritten as
$\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}P_{k}^{T}B_{k}P_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+g_{k}^{T}P_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix},$
having set $P_{k}=\begin{bmatrix}-g_{k}&s_{k}\end{bmatrix}$.
Then, proposition 3 shows that condition eq. 14 on the eigenvalues of the matrices $B_{k}$ guarantees that the solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ of (8) produces directions $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$ that are gradient-related. However, checking or ensuring that a sequence of $n\times n$ matrices $\{B_{k}\}$ have uniformly bounded eigenvalues can be extremely challenging for large-dimensional optimization problems.
Problem (8) can be rewritten in the following form:
$\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}H_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (22)
In this section we analyze conditions on the $2\times 2$ matrix $H_{k}$ sufficient to guarantee suitable properties to the direction $d_{k}$.
This analysis is fundamental for two reasons. The first one is that, as said before, these “direct” conditions on the low-dimensional matrix can be checked and ensured (possibly by suitable modifications) regardless of the $n\times n$ matrix $B_{k}$. The second reason is that they can be employed in connection with any $2\times 2$ matrix defining a quadratic model — see (11) — to be minimized for determining the values $\alpha_{k}$ and $\beta_{k}$ that characterize the update rule of gradient methods with momentum.
Following the latter point, we assume that $H_{k}$ is any symmetric positive definite matrix. Then, (6) is a particular case of $H_{k}$ provided that $B_{k}$ is positive definite and that $g_{k}$ and $s_{k}$ are linearly independent.
A possible approach to define conditions to impose on the sequence of matrices $\{H_{k}\}$ might, in principle, draw inspiration from proposition 3, concerning the sequence of matrices $\{B_{k}\}$. First, we therefore state the following theorem involving the sequence of matrices $\{H_{k}\}$.
Proposition 4.
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k},$ where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is the solution of (22).
Let $\{H_{k}\}\subseteq\mathbb{R}^{2\times 2}$ be the sequence of symmetric matrices defining problems (22) and assume that there exist scalars $0<\hat{c}_{1}\leq\hat{c}_{2}$ such that
$\hat{c}_{1}\leq\lambda_{\min}(H_{k})\leq\lambda_{\max}(H_{k})\leq\hat{c}_{2}$ (23)
holds for all $k$.
Then the direction $d_{k}$ satisfies the following conditions:
$g_{k}^{T}d_{k}\leq-\frac{1}{\hat{c}_{2}}\|g_{k}\|^{4}$ (24)
$\frac{1}{\hat{c}_{2}}\|g_{k}\|^{3}\leq\|d_{k}\|\leq\frac{1}{\hat{c}_{1}}(\|g_{k}\|+\|s_{k}\|)(\|g_{k}\|^{2}+|g_{k}^{T}s_{k}|).$ (25)
Proof.
We know that $\alpha_{k}$ and $\beta_{k}$ are such that:
$\begin{bmatrix}\alpha_{k}\\ \beta_{k}\end{bmatrix}\ =\ {H_{k}}^{-1}\begin{bmatrix}\|g_{k}\|^{2}\\ -g_{k}^{T}s_{k}\end{bmatrix}.$ (26)
Multiplying both sides of the above equation by $\begin{bmatrix}\|g_{k}\|^{2}&-g_{k}^{T}s_{k}\end{bmatrix}$ we obtain:
$\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}=\begin{bmatrix}\|g_{k}\|^{2}&-g_{k}^{T}s_{k}\end{bmatrix}H_{k}^{-1}\begin{bmatrix}\|g_{k}\|^{2}\\ -g_{k}^{T}s_{k}\end{bmatrix}.$ (27)
We note that $-g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$; thus, by using eq. 23 and (27) we have:
$-g_{k}^{T}d_{k}\geq\lambda_{min}(H^{-1}_{k})\biggl{(}\|g_{k}\|^{4}+(g_{k}^{T}s_{k})^{2}\biggr{)},$
which implies:
$g_{k}^{T}d_{k}\leq-\frac{1}{\hat{c}_{2}}\|g_{k}\|^{4},$
and thus proves (24).
The previous inequality and the Schwarz inequality imply
$\|d_{k}\|\geq\frac{1}{\hat{c}_{2}}\|g_{k}\|^{3}$
which proves the first inequality of (25).
By using again (26), we can get an upper bound on the norm of $d_{k}$:
$\|d_{k}\|=\biggl{\|}\begin{bmatrix}-g_{k}&s_{k}\end{bmatrix}\begin{bmatrix}\alpha_{k}\\ \beta_{k}\end{bmatrix}\ \biggr{\|}=\biggl{\|}\ \begin{bmatrix}-g_{k}&s_{k}\end{bmatrix}{H_{k}}^{-1}\begin{bmatrix}\|g_{k}\|^{2}\\ -g_{k}^{T}s_{k}\end{bmatrix}\biggr{\|}$
$\qquad\ \leq\biggl{\|}{H_{k}}^{-1}\biggr{\|}\ \biggl{\|}\begin{bmatrix}-g_{k}&s_{k}\end{bmatrix}\biggr{\|}\ \biggl{\|}\begin{bmatrix}\|g_{k}\|^{2}\\ -g_{k}^{T}s_{k}\end{bmatrix}\biggr{\|}$
$\qquad\ \leq\frac{1}{\hat{c}_{1}}(\|g_{k}\|+\|s_{k}\|)(\|g_{k}\|^{2}+|g_{k}^{T}s_{k}|),$
which proves the second relation of (25) and completes the proof. ∎
According to the above result, suitable descent properties of the obtained search directions hold and would in fact be sufficient, coupled with the employment of a standard Armjio-type line search, to define a globally convergent gradient method with momentum.
However, as we detail below, the obtained sequence of directions $\{d_{k}\}$ is not gradient-related according to definition 1 and, hence, an ${\cal O}(\epsilon^{-2})$ complexity bound cannot be ensured.
Remark 1.
We observe that when the matrix $H_{k}$ is given by (6), it tends to become the null matrix as $x_{k}$ approaches a stationary point. Hence, uniform boundedness conditions on the eigenvalues of $H_{k}$ are in contrast with the matrices deriving from (3).
Even more in general, under the assumptions of proposition 4, it is actually not possible to ensure that the obtained sequence of directions $\{d_{k}\}$ is gradient-related. Indeed, consider the cases where $g_{k}^{T}s_{k}=0$. The inequality (25) implies:
$|g_{k}^{T}d_{k}|\leq\|g_{k}\|\|d_{k}\|\leq\frac{1}{\hat{c}_{1}}(\|g_{k}\|+\|s_{k}\|)\|g_{k}\|^{3},$
and hence, by assuming that $\{s_{k}\}$ is bounded, for sufficiently small values of $\|g_{k}\|$, the direction $d_{k}$ does not satisfy the first requirement of (13).
By the next proposition, we finally state suitable conditions on the sequence of matrices $\{H_{k}\}$, that also take into account the sequences of vectors involved, i.e., $\{g_{k}\}$ and $\{s_{k}\}$. These conditions are sufficient to ensure that the obtained sequence of search directions $\{d_{k}\}$ is gradient-related.
Proposition 5.
Let $\{H_{k}\}\subseteq\mathbb{R}^{2\times 2}$ be the sequence of symmetric matrices defining problems (22) and assume that there exist scalars $0<c_{1}\leq c_{2}$ such that
$c_{1}\leq\lambda_{\min}(D_{k}^{-1}H_{k}D_{k}^{-1})\leq\lambda_{\max}(D_{k}^{-1}H_{k}D_{k}^{-1})\leq c_{2}$ (28)
holds for all $k$, where
$D_{k}=\begin{bmatrix}\|g_{k}\|&0\\ 0&\|s_{k}\|\end{bmatrix}.$
Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of problem (22). Then, the direction $d_{k}$ satisfies the following conditions:
$g_{k}^{T}d_{k}\leq-\frac{1}{c_{2}}\|g_{k}\|^{2}$ (29)
$\frac{1}{c_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{c_{1}}\|g_{k}\|.$ (30)
Proof.
For simplicity, let us introduce the following matrix:
$\tilde{H}_{k}=D_{k}^{-1}H_{k}D_{k}^{-1}=\begin{bmatrix}\frac{(H_{11})_{k}}{\|g_{k}\|^{2}}&\frac{(H_{12})_{k}}{\|s_{k}\|\,\|g_{k}\|}\\ \frac{(H_{12})_{k}}{\|s_{k}\|\,\|g_{k}\|}&\frac{(H_{22})_{k}}{\|s_{k}\|^{2}}\end{bmatrix}$
Then, problem (22) can be rewritten as:
$\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}D_{k}\tilde{H}_{k}D_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix},$
and its optimal solutions $\alpha_{k}$ and $\beta_{k}$ are such that
$D_{k}\tilde{H}_{k}D_{k}\begin{bmatrix}\alpha_{k}\\ \beta_{k}\end{bmatrix}\ =\ \begin{bmatrix}\|g_{k}\|^{2}\\ -g_{k}^{T}s_{k}\end{bmatrix},$ (31)
from which we have:
$D_{k}\begin{bmatrix}\alpha_{k}\\ \beta_{k}\end{bmatrix}\ =\ {\tilde{H}_{k}}^{-1}D_{k}^{-1}\begin{bmatrix}\|g_{k}\|^{2}\\ -g_{k}^{T}s_{k}\end{bmatrix},$
namely:
$\begin{bmatrix}\alpha_{k}\|g_{k}\|\\ \beta_{k}\|s_{k}\|\end{bmatrix}\ =\ {\tilde{H}_{k}}^{-1}\begin{bmatrix}\|g_{k}\|\\ -\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}.$
Multiplying both sides of the above equation by $\begin{bmatrix}\|g_{k}\|&-\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}$ we obtain:
$\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}=\begin{bmatrix}\|g_{k}\|&-\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}\tilde{H}_{k}^{-1}\begin{bmatrix}\|g_{k}\|\\ -\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}.$ (32)
Recalling eq. 28 and that $-g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$, we have from (32):
$-g_{k}^{T}d_{k}\geq\lambda_{min}(\tilde{H}^{-1})\biggl{(}\|g_{k}\|^{2}+\left(\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\right)^{2}\biggr{)},$
and thus
$g_{k}^{T}d_{k}\leq-\frac{1}{c_{2}}\|g_{k}\|^{2}.$ (33)
Then, Schwarz inequality implies:
$\|g_{k}\|\leq c_{2}\,\|d_{k}\|.$ (34)
Now, multiplying both terms of equality (31) by $\begin{bmatrix}\alpha_{k}&\beta_{k}\end{bmatrix}$ we obtain:
$\begin{bmatrix}\alpha_{k}\|g_{k}\|\\ \beta_{k}\|s_{k}\|\end{bmatrix}^{T}\tilde{H}_{k}\begin{bmatrix}\alpha_{k}\|g_{k}\|\\ \beta_{k}\|s_{k}\|\end{bmatrix}\ =-g_{k}^{T}(-\alpha g_{k}+\beta_{k}s_{k})=-g_{k}^{T}d_{k},$
from which we get:
$-g_{k}^{T}d_{k}$ $\geq\lambda_{min}\left(\tilde{H}_{k}\right)(\alpha_{k}^{2}\|g_{k}\|^{2}+\beta_{k}^{2}\|s_{k}\|^{2})$
$\geq c_{1}(\alpha_{k}^{2}\|g_{k}\|^{2}+\beta_{k}^{2}\|s_{k}\|^{2})$
$\geq\frac{c_{1}}{2}\|d_{k}\|^{2}.$
Then, by using Schwarz inequality we get:
$\|g_{k}\|\|d_{k}\|\geq|g_{k}^{T}d_{k}|\geq\frac{c_{1}}{2}\|d_{k}\|^{2}$
and hence
$\|d_{k}\|\leq\frac{2}{c_{1}}\|g_{k}\|.$ (35)
Now (33), (34) and (35) imply (29) and (30) and, hence, the proof is complete. ∎
Remark 2.
It is important to note that assumption (28) is not difficult to satisfy. In fact, given any sequence of symmetric matrices $\{\hat{H}_{k}\}\subseteq\mathbb{R}^{2\times 2}$ for which there exist scalars $0<\tilde{c}_{1}\leq\tilde{c}_{2}$ such that, for all $k$,
$\tilde{c}_{1}\leq\lambda_{\min}(\hat{H}_{k})\leq\lambda_{\max}(\hat{H}_{k})\leq\tilde{c}_{2},$ (36)
a sequence of matrices $\{H_{k}\}$ satisfying (28) is obtained by choosing:
$H_{k}=\left[\begin{array}[]{cc}\|g_{k}\|(\hat{H}_{11})_{k}\|g_{k}\|&\|g_{k}\|(\hat{H}_{12})_{k}\|s_{k}\|\\ \|s_{k}\|(\hat{H}_{21})_{k}\|g_{k}\|&\|s_{k}\|(\hat{H}_{22})_{k}\|s_{k}\|\end{array}\right].$
5 Algorithmic model
In this section we propose an algorithmic framework for gradient methods with momentum which exploits the theoretical analysis carried out in the previous sections. The idea is to define a first-order algorithm with both strong theoretical guarantees in the nonconvex setting and the possibility of computationally exploiting eventual second-order information on the minimization problem. Before formally presenting the algorithm, we summarize the key steps:
  • (a)
    define a quadratic subproblem
    $\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}H_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (37)
    where $H_{k}$ is a $2\times 2$ symmetric matrix;
  • (b)
    once computed a solution $[\alpha_{k}\ \beta_{k}]^{T}$ of (37), provided one exists, a check on the obtained search direction
    $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$
    is performed in order to ensure the gradient-related property of the sequence $\{d_{k}\}$;
  • (c)
    if the test is satisfied, then a standard Armjio-type line search is performed along $d_{k}$; otherwise, a suitable modification of $H_{k}$ based on (28) is introduced and, again, steps (a) (with the modified $H_{k}$) and (b) (without the check on $d_{k}$) are performed, as well as the Armjio-type line search along the obtained $d_{k}$.
The proposed Algorithmic Model is described in Algorithm 1. Notice that the initial tentative stepsize $\eta=1$ is optimal according to the quadratic model used to define the search direction $d_{k}$; thus, the unit step will often be a good step even for the true objective and satisfy the Armijo sufficient decrease condition. From a computational perspective, this allows to possibly save several backtracking steps and, consequently, function evaluations.
1 Input: $x_{0}\in\mathbb{R}^{n}$, $\gamma\in(0,1)$, $\delta\in(0,1)$, $c_{1}>0$, $c_{2}>0$.
2 Set $k\leftarrow 0$
3 while $\|\nabla f(x_{k})\|>0$ do
       /* Compute the search direction */
4       Set gr_dir_found $\leftarrow$ False
5       Define a $2\times 2$ symmetric matrix $H_{k}$
6       if problem (37) admits solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ then
7             Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
8              if $g_{k}^{T}d_{k}\leq-c_{1}\|g_{k}\|^{2}$ and $\|d_{k}\|\leq c_{2}\|g_{k}\|$  then
9                    Set gr_dir_found $\leftarrow$ True
10
11      if gr_dir_found $=$ False then
12             Define a new $2\times 2$ symmetric matrix $H_{k}$ satisfying condition (28)
13              Compute $\alpha_{k}$ and $\beta_{k}$ by solving (37)
14             Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
15
      /* Perform Armijo line search along $d_{k}$/
16
17      Set $\eta\leftarrow 1$
18       while $f(x_{k}+\eta d_{k})>f(x_{k})+\gamma\eta d_{k}^{\top}\nabla f(x_{k})$ do
19              Set $\eta\leftarrow\delta\eta$
20
21      Set $\eta_{k}\leftarrow\eta$, $x_{k+1}\leftarrow x_{k}+\eta_{k}d_{k}$
22       Set $k\leftarrow k+1$
Algorithm 1 Gradient Method with Momentum (GMM)
The theoretical properties of the proposed framework for gradient methods with momentum are stated in the following proposition.
Proposition 6.
Assume that $\mathcal{L}_{0}=\{x\in\mathbb{R}^{n}\mid f(x)\leq f(x_{0})\}$ is a compact set. Then, Algorithm 1 either stops in a finite number of iterations $\nu$ producing a point $x_{\nu}$ which is stationary for $f$, i.e. $\nabla f(x_{\nu})=0$, or it produces an infinite sequence $\{x_{k}\}$ that admits limit points, each one being a stationary point for $f$. Furthermore, if the gradient $\nabla f$ is Lipschitz continuous on $\mathbb{R}^{n}$, we have that Algorithm 1 requires at most $\mathcal{O}(\epsilon^{-2})$ iterations, function and gradient evaluations to attain
$\|\nabla f(x_{k})\|\leq\epsilon_{k}.$
Proof.
From the steps the algorithm and Proposition 5 we have that $\{d_{k}\}$ is a sequence of gradient-related directions. Since the Armijo line search is employed within Algorithm 1, the results follow from [24, Proposition 1.2.1] and by [23]. ∎
We now focus on the issue of modifying a given $2\times 2$ matrix $H_{k}$ in order to satisfy condition (28) as required at step 1 of Algorithm 1. Suppose that a symmetric $H_{k}$ matrix has been defined at step 1, but either problem problem (37) does not admit solution or the test at step 1 is not satisfied, i.e., the value of gr_dir_found remains False. Then, step 1 must be performed and a new matrix $H_{k}$ must be constructed modifying as least as possible the given matrix defined at step 1.
Let us denote by $H_{k}^{0}$ the matrix defined at step 1 and by $H_{k}$ the new matrix defined at step 1. We can proceed as follows:
    Let $\hat{H}_{k}$ be obtained by a modified Cholesky factorization [24] applied to ${D_{k}^{-1}}H_{k}^{0}D_{k}^{-1}$, where
    $D_{k}=\begin{bmatrix}\|g_{k}\|&0\\ 0&\|s_{k}\|\end{bmatrix};$
    Set $H_{k}=D_{k}\hat{H}_{k}D_{k}$, i.e.,
    $H_{k}=\left[\begin{array}[]{cc}\|g_{k}\|(\hat{H}_{11})_{k}\|g_{k}\|&\|g_{k}\|(\hat{H}_{12})_{k}\|s_{k}\|\\ \|s_{k}\|(\hat{H}_{21})_{k}\|g_{k}\|&\|s_{k}\|(\hat{H}_{22})_{k}\|s_{k}\|\end{array}\right].$
The boundedness of $\{H_{k}^{0}\}$ and the properties of the modified Cholesky factorization imply that (36) of Remark 2 holds, so that, according to the same remark, the matrix $H_{k}$ is such that condition (28) is satisfied.
We conclude this section by stating a theoretical result showing that, under suitable assumptions on the objective function and a choice of the matrix $H_{k}$ actually related to the Hessian $\nabla^{2}f(x_{k})$, steps 1-1 are never executed for $k$ sufficiently large, that is, the test at step 1 is always satisfied.
Proposition 7.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a twice continuously differentiable function. Suppose that
$H_{k}=\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}$
is the matrix defined at step 1, and let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Assume that $\{x_{k}\}$ converges to $x^{*}$, where $\nabla f(x^{*})=0$ and the Hessian matrix $\nabla^{2}f(x^{*})$ is positive definite. Furthermore, assume that the constants of the test at step 1 are chosen in a such a way that:
$c_{1}\leq\theta^{3}\frac{\lambda_{\min}(\nabla^{2}f(x^{*}))}{\lambda_{\max}(\nabla^{2}f(x^{*}))^{2}},\qquad\qquad c_{2}\geq\frac{2}{\theta\lambda_{\min}(\nabla^{2}f(x^{*}))}.$
where $\theta\in(0,1)$. Then, for $k$ sufficiently large the test at step 1 is satisfied.
Proof.
By continuity there exists a neighborhood ${\cal B}(x^{*})$ of $x^{*}$ such that if $x_{k}\in{\cal B}(x^{*})$ we have:
$\theta\lambda_{\min}(\nabla^{2}f(x^{*}))\leq\lambda_{\min}(\nabla^{2}f(x_{k})),\qquad\lambda_{\max}(\nabla^{2}f(x_{k}))\leq\frac{1}{\theta}\lambda_{\max}(\nabla^{2}f(x^{*})).$
Then the thesis follows from Proposition 3 by setting
$B_{k}=\nabla^{2}f(x_{k}),\quad\eta_{1}=\theta\lambda_{\min}(\nabla^{2}f(x_{k})),\quad\eta_{2}=\frac{1}{\theta}\lambda_{\max}(\nabla^{2}f(x_{k}))$
and the proof is complete. ∎
6 Computation of ${H_{k}}$
The core of the general framework lies in how a sequence of $2\times 2$ matrices $\{H_{k}\}$ can be determined to ensure an efficient computational behavior of the algorithm, as well as exploiting proposition 3 or proposition 5 to guarantee sound theoretical properties.
Regarding the first issue, the following two strategies can be adopted:
  • i)
    define a suitable $n\times n$ matrix $B_{k}$ and compute the matrix $H_{k}$ by using (6);
  • ii)
    define a suitable $2\times 2$ matrix $H_{k}$ independent on any $n\times n$ matrix.
Concerning strategy i), in large-scale optimization problems the use and storage of the $n\times n$ matrix $B_{k}$ can be computationally too expensive, if not downright prohibitive. Therefore, to take into account this issue, we propose two approaches described in section 6.1 and section 6.3. A technique related to strategy ii) is presented in section 6.2. Summarizing, we propose three techniques to compute the matrix $H_{k}$, although other approaches could be exploited within the general framework we have presented.
6.1 Approximating Hessian-vector products by finite differences of gradients
In line with the strategy employed in [20, 21], we can draw inspiration by the Truncated Newton methods approach [25, 26], where the explicit management of the Hessian matrix $\nabla^{2}f(x_{k})$ is not required, but rather the Hessian-vector product $\nabla^{2}f(x_{k})d$ is directly handled, with $d\in R^{n}$.
Consider $H_{k}$ defined by (6), i.e.,
$H_{k}=\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix},$
The elements of $H_{k}$ can be obtained estimating the two matrix-vector products $B_{k}g_{k}$ and $B_{k}s_{k}$ by finite difference approximation, namely by the following vectors (where $\xi>0$ is a suitably small parameter):
$\frac{\nabla f(x_{k}+\xi g_{k}/\|g_{k}\|)-g_{k}}{\xi/\|g_{k}\|}$$\approx$ $\nabla^{2}f(x_{k})g_{k},$
$\frac{\nabla f(x_{k}+\xi s_{k}/\|s_{k}\|)-g_{k}}{\xi/\|s_{k}\|}$$\approx$ $\nabla^{2}f(x_{k})s_{k}.$
In this way it is possible to consistently construct $H_{k}$ without the need of handling an $n\times n$ matrix. The price to pay consists in two additional evaluations of the $n$ dimensional gradient of $f$, $\nabla f(x)$.
6.2 Hessian estimation in subspace by interpolation
In this subsection, an alternative way to compute the matrix $H_{k}$ is proposed that avoids the need of additional $n$ dimensional gradient evaluations. We note that a simplified version of the following approach was found to be particularly well-performing in [20].
Let us consider the two-dimensional function
$\psi_{k}(\alpha,\beta)=f(x_{k}-\alpha g_{k}+\beta s_{k}).$
Assuming that $f$ is twice continuously differentiable, we have seen — see (10) — that the quadratic Taylor polynomial of $\psi_{k}(\alpha,\beta)$ centered at $(0,0)$ is
$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$
Then, the matrix $H_{k}$ defined by (6) can be viewed as a matrix approximating $\nabla^{2}\psi_{k}(0,0)$, and, in a more general setting, $H_{k}$ could be any $2\times 2$ matrix (independent of the $n\times n$ matrix $B_{k}$). This leads to consider the following approximation of the quadratic Taylor polynomial of $\psi_{k}(\alpha,\beta)=f(x_{k}-\alpha g_{k}+\beta s_{k})$:
$\phi(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\left[\begin{array}[]{cc}H_{11}&H_{12}\\ H_{12}&H_{22}\end{array}\right]\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$
The three elements defining a symmetric matrix $H_{k}$ can be determined by imposing the interpolation conditions on three points $(\alpha_{1},\beta_{1})$, $(\alpha_{2},\beta_{2})$, and $(\alpha_{3},\beta_{3})$ different from $(0,0)$:
$\phi(\alpha_{1},\beta_{1})=\psi_{k}(\alpha_{1},\beta_{1})=f(x_{k}-\alpha_{1}g_{k}+\beta_{1}s_{k}),$
$\phi(\alpha_{2},\beta_{2})=\psi_{k}(\alpha_{2},\beta_{2})=f(x_{k}-\alpha_{2}g_{k}+\beta_{2}s_{k}),$
$\phi(\alpha_{3},\beta_{3})=\psi_{k}(\alpha_{3},\beta_{3})=f(x_{k}-\alpha_{3}g_{k}+\beta_{3}s_{k}),$
i.e., by solving the linear system
$\frac{1}{2}\left(\!\!\begin{array}[]{ccc}\alpha_{1}^{2}&2\alpha_{1}\beta_{1}&\beta_{1}^{2}\\ \alpha_{2}^{2}&2\alpha_{2}\beta_{2}&\beta_{2}^{2}\\ \alpha_{3}^{2}&2\alpha_{3}\beta_{3}&\beta_{3}^{2}\end{array}\!\!\right)\!\!\left(\!\!\begin{array}[]{c}H_{11}\\ H_{12}\\ H_{22}\end{array}\!\!\right)=$
$\qquad\qquad\left(\!\!\begin{array}[]{c}f(x_{k}-\alpha_{1}g_{k}+\beta_{1}s_{k})-f(x_{k})+\alpha_{1}\|g_{k}\|^{2}-\beta_{1}g_{k}^{T}s_{k}\\ f(x_{k}-\alpha_{2}g_{k}+\beta_{2}s_{k})-f(x_{k})+\alpha_{2}\|g_{k}\|^{2}-\beta_{2}g_{k}^{T}s_{k}\\ f(x_{k}-\alpha_{3}g_{k}+\beta_{3}s_{k})-f(x_{k})+\alpha_{3}\|g_{k}\|^{2}-\beta_{3}g_{k}^{T}s_{k}\end{array}\!\!\right).$
Remark 3.
The described strategy requires three function evaluations. However, considering the point $(\alpha_{1},\beta_{1})=(0,-1)$, we have
$f(x_{k}-\alpha_{1}g_{k}+\beta_{1}s_{k})=f(x_{k}-(x_{k}-x_{k-1})=f(x_{k-1})$
Then, by exploiting the information of the past iteration, i.e., the knowledge of $f(x_{k-1})$, the matrix $H_{k}$ can be built by only two additional function evaluations.
Two reasonable candidate points to evaluate the function at might be, for example, $(\alpha_{k-1},\beta_{k-1})$ and $(\alpha_{k-1},0)$. The interpolation system to obtain the quadratic matrix $H_{k}$ becomes
$\frac{1}{2}\left(\begin{array}[]{ccc}0&0&1\\ \alpha_{k-1}^{2}&0&0\\ \alpha_{k-1}^{2}&2\alpha_{k-1}\beta_{k-1}&\beta_{k-1}^{2}\end{array}\right)\left(\begin{array}[]{c}H_{11}\\ H_{12}\\ H_{22}\end{array}\right)=$
$\qquad\qquad\left(\begin{array}[]{c}f(x_{k-1})-f(x_{k})+g_{k}^{T}s_{k}\\ f(x_{k}-\alpha_{k-1}g_{k})-f(x^{k})+\alpha_{k-1}\|g_{k}\|^{2}\\ f(x_{k}-\alpha_{k-1}g_{k}+\beta_{k-1}s_{k})-f(x_{k})+\alpha_{k-1}\|g_{k}\|^{2}-\beta_{k-1}g_{k}^{T}s_{k}\end{array}\right).$
6.3 Hessian approximation by a diagonal matrix
The last example of computation of the matrix $H_{k}$ does not require any additional function or gradient computations. The idea is to consider a diagonal matrix $B_{k}$, that is
$B_{k}=\begin{bmatrix}(\mu_{k})_{1}&0&\cdots&0\\ 0&(\mu_{k})_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&\cdots&\cdots&(\mu_{k})_{n}\end{bmatrix}.$ (42)
The diagonal elements of the matrix $B_{k}$ are computed by drawing inspiration from the approach of Barzilai-Borwein methods [27]. The matrix $B_{k}$ given by (42) is the optimal solution of the following problem:
$\min_{B}\|B_{k}s_{k}-y_{k}\|^{2}$
where
$y_{k}=g_{k}-g_{k-1},\qquad s_{k}=x_{k}-x_{k-1}.$
This implies that, for $i=1,\ldots,n$:
$(\mu_{k})_{i}=(y_{k})_{i}/(s_{k})_{i}.$
Finally, the elements of the $2\times 2$ matrix $H_{k}$ are given by:
$(H_{11})_{k}=\sum_{i=1}^{n}(\mu_{k})_{i}(g_{k})_{i}^{2},\quad(H_{12})_{k}=\sum_{i=1}^{n}(\mu_{k})_{i}(g_{k})_{i}(s_{k})_{i},\quad(H_{22})_{k}=\sum_{i=1}^{n}(\mu_{k})_{i}(s_{k})_{i}^{2}.$
7 Computational experiments
In this section, we describe and report the results of thorough computational experiments aimed at assessing the potential of the algorithm proposed in this work. Specifically, we considered to this aim a benchmark of 163 problems from the CUTEst test-suite [28]. In particular, we included in our test set all 183 problems in the CUTEst collection that are 1) unconstrained, 2) regular, 3) with a number of variables $n$ that is greater or equal than 1000 (except for the problems LUKSAN12LS, LUKSAN13LS, LUKSAN14LS having 998 variables) or that is user definable; in the latter case, we set the number of variables to 1000 or as close as possible to 1000. In Table 1 those 183 problems are reported along with their dimensions.
Table 1: Collection of 183 unconstrained CUTEst problems ($n$ denotes the number of variables). The problems whose name is crossed out have then been removed from the initial selection as they turned out to be unbounded from below.
Problem $n$ Problem $n$ Problem $n$
ARGLINA 1000 EDENSCH 2000 OBSTCLBU 9604
ARGLINB 1000 EIGENALS 2550 ODC 4900
ARGLINC 1000 EIGENBLS 2550 ODNAMUR 11130
ARGTRIGLS 1000 EIGENCLS 2652 OSCIGRAD 100000
BDEXP 5000 ENGVAL1 5000 OSCIPATH 5000
BOX 10000 EXTROSNB 1000 PENALTY1 1000
BOXPOWER 10000 FLETBV3M 5000 PENALTY2 1000
BRATU1D 5001 FLETCBV2 5000 PENALTY3 1000
BROWNAL 1000 FLETCBV3 5000 PENTDI 5000
BROYDN3DLS 5000 FLETCHBV 5000 POWELLBC 1000
BROYDNBDLS 5000 FLETCHCR 1000 POWELLSG 5000
CHEBYQAD 1000 FMINSRF2 5625 POWER 10000
CHNRSNBM 1000 FMINSURF 5625 POWERSUM 1000
CLPLATEA 4970 FREUROTH 5000 PRICE3 2000
CURLY30 10000 GENHUMPS 5000 PROBPENL 1000
CVXBQP1 10000 GENROSE 1000 QING 1000
CYCLIC3LS 100002 GENROSEB 1000 QRTQUAD 5000
CYCLOOCFLS 29996 GRIDGENA 5560 QUARTC 5000
CYCLOOCTLS 29996 HILBERTA 1000 RAYBENDL 2046
DEGDIAG 100001 HILBERTB 1000 RAYBENDS 2046
DEGTRID 100001 INDEF 5000 S368 1000
ARWHEAD 5000 INDEFM 100000 SBRYBND 5000
BDQRTIC 5000 INTEQNELS 1002 SCHMVETT 5000
BROYDN7D 5000 JNLBRNG1 9604 SCOSINE 5000
BRYBND 5000 JNLBRNG2 9604 SCURLY10 10000
CHAINWOO 4000 JNLBRNGA 9604 SCURLY20 10000
CLPLATEB 4970 JNLBRNGB 9604 SCURLY30 10000
CLPLATEC 4970 KSSLS 1000 SENSORS 1000
COSINE 10000 LIARWHD 5000 SINEALI 1000
CRAGGLVY 5000 LINVERSE 1999 SINQUAD 5000
CURLY10 10000 LMINSURF 5329 SPARSINE 5000
CURLY20 10000 LUKSAN11LS 1000 SPARSQUR 10000
DEGTRID2 100001 LUKSAN12LS 998 SPIN2LS 1002
DIAGIQB 1000 LUKSAN13LS 998 SPINLS 1327
DIAGIQE 1000 LUKSAN14LS 998 SROSENBR 5000
DIAGIQT 1000 LUKSAN15LS 1000 SSBRYBND 5000
DIAGNQB 1000 LUKSAN16LS 1000 SSC 4900
DIAGNQE 1000 LUKSAN17LS 1000 SSCOSINE 5000
DIAGNQT 1000 LUKSAN21LS 1000 STRTCHDV 1000
DIAGPQB 1000 LUKSAN22LS 1000 TESTQUAD 5000
DIAGPQE 1000 MANCINO 1000 TOINTGSS 5000
DIAGPQT 1000 MCCORMCK 5000 TORSION1 5184
DIXMAANA1 3000 MODBEALE 20000 TORSION2 5184
DIXMAANB 3000 MOREBV 5000 TORSION3 5184
DIXMAANC 3000 NCB20 5010 TORSION4 5184
DIXMAAND 3000 NCB20B 5000 TORSION5 5184
DIXMAANE1 3000 NCVXBQP1 10000 TORSION6 5184
DIXMAANF 3000 NCVXBQP2 10000 TORSIONA 5184
DIXMAANG 3000 NCVXBQP3 10000 TORSIONB 5184
DIXMAANH 3000 NLMSURF 5329 TORSIONC 5184
DIXMAANI1 3000 NOBNDTOR 5184 TORSIOND 5184
DIXMAANJ 3000 NONCVXU2 5000 TORSIONE 5184
DIXMAANK 3000 NONCVXUN 5000 TORSIONF 5184
DIXMAANL 3000 NONDIA 5000 TQUARTIC 5000
DIXMAANM1 3000 NONDQUAR 5000 TRIDIA 5000
DIXMAANN 3000 NONMSQRT 4900 TRIGON1 1000
DIXMAANO 3000 NONSCOMP 5000 TRIGON2 1000
DIXMAANP 3000 OBSTCLAE 9604 VANDANMSLS 1002
DIXON3DQ 10000 OBSTCLAL 9604 VARDIM 1000
DQDRTIC 5000 OBSTCLBL 9604 VAREIGVL 5000
DQRTIC 5000 OBSTCLBM 9604 WOODS 4000
Then, from this set of 183 problems, we removed those problems that are unbounded below, i.e., the problems where at least one solver among the ones considered generates a sequence $\{x_{k}\}$ such that $f(x_{k})\to-\infty$. Those problems are marked in table 1 with a strike-through.
The code for all the experiments described in this work was written in Python 3.9, only exploiting numpy and scipy libraries. The software is available at:
Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rules discussed in section 6; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq. 28 holds. As for the parameters, we set $\delta=0.5,\;\gamma=10^{-5}$.
The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration. As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-3}$; we also set a maximum number of iterations to 5000, so that we consider a failure each run that ends by hitting this threshold.
As a baseline for comparisons, we took into account classical nonlinear conjugate gradient methods [29] and the L-BFGS algorithm [30], which can arguably be considered the state-of-the-art for the solution of smooth unconstrained nonlinear optimization problems. For both algorithms, we considered the very efficient implementations available through the scipy library.
The results of the experiments are shown in the form of performance profiles [31]. As performance metrics, we considered both the runtime and the number of iterations.
In the following, we report results for three versions of our algorithm, depending on the technique adopted for the computation of $H_{k}$, as detailed in section 6. More in detail, we denote by GMM1, GMM2, GMM3 the algorithm where $H_{k}$ is computed as described in subsection 6.1, 6.2 and 6.3, respectively.
Figure 1: Performance profiles on all 163 problems.
Figure 2: Performance profiles on the subset of problems where all the solvers find the same best function value.
In fig. 1, we report the performance profiles in terms of time and number of iterations for the comparison among GMM1, GMM2, GMM3 and CGscipy on the whole collection of 163 problems. As we can see, the best performing solver is GMM2 both in terms of time and iterations. We can also say that GMM3 is the second best solver in terms of time even though it is the worst one in terms of iterations. Both solvers GMM2 and GMM3 are better than CGscipy in terms of time. In doing these considerations, we must also take into account that it might happen that solvers converge toward different points. Indeed, in table 2 we report the number of wins, i.e., the number of problems where each solver finds the best function value.
Table 2: Number of wins, i.e., the number of problems where each solver finds the best function value, on all 163 test problems. GMM variants and CGscipy solvers are considered.
Solver # wins
GMM1 71
GMM2 58
GMM3 26
CGscipy 16
Then, in fig. 2, we report the performance profiles considering the subset of problems where each solver finds the same best function value, i.e.
$f^{*}_{i}-f_{L}<10^{-3}\quad\forall\ i\in\cal S,$
where $\cal S$ is the set of solvers. From fig. 2, we note that the performance of GMM1 becomes closer to that of the conjugate gradient method in terms of computational time, whereas GMM2 and GMM3 continue to be clearly superior.
Figure 3: Performance profiles for the comparison between GMM1, GMM2, GMM3 and L-BFGS on all the 163 problems.
Figure 4: Performance profiles for the comparison between GMM1, GMM2, GMM3 and L-BFGS on the subset of problems where all the solvers find the same best function value.
Now, we consider the comparison between our methods and the state-of-the-art algorithm for unconstrained optimization, namely, L-BFGS [30]. In particular, performance profiles of solvers GMM1, GMM2, GMM3 and L-BFGS on the entire collection of 163 problems are reported in fig. 3. As we can see, our method GMM2 is quite competitive with L-BFGS in terms of computational time. Again, we have to consider the fact that the different solvers can converge toward points with different function values. In fact, as reported in table 3, GMM1 and GMM2 find a better function value more frequently than L-BFGS.
Table 3: Number of wins, i.e., the number of problems where each solver finds the best function value, on all 163 test problems. $GMM$ variants and L-BFGS solvers are considered.
Solver # wins
GMM1 55
GMM2 40
GMM3 17
L-BFGSscipy 33
In fig. 4, we thus consider the performance profiles obtained for the four solvers on the subset of problems where all of them find the same function value. As we can see, GMM2 gained some efficiency and it now slightly dominates L-BFGS in terms of computational time.
Then, for major clarity, we report in figs. 5 and 6 the performance profiles for the comparison between our bet method GMM2 and L-BFGS.
Figure 5: Performance profiles for the comparison between GMM2 and L-BFGS on all the 163 problems.
Figure 6: Performance profiles for the comparison between GMM2 and L-BFGS on the subset of problems where all the solvers find the same best function value.
From figs. 5 and 6, we can say that GMM2 is quite competitive if not better than L-BFGS both in terms of computational time and number of iterations. The advantages of using GMM2 are more evident when the two solvers are compared on the subset of problems where they both find the same function value.
8 Conclusions
In this work, we introduced a general framework of gradient methods with momentum for nonconvex optimization. For the proposed class of algorithms, we proved global convergence and optimal worst-case complexity bounds. The assumptions required to obtain the theoretical results are lighter to check, from a computational perspective, than those required in related works from the literature. This result allowed to devise particularly efficient ways of implementing the proposed method. Thorough computational results we showed that the novel algorithm outperforms standard conjugate gradient methods and is competitive with L-BFGS, the state-of-the-art for nonlinear optimization problems.
Future studies might concern improvements of the proposed framework and suitable modifications to make it exploitable in different settings.
Declarations
Funding
No funding was received for conducting this study.
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Code Availability Statement
The code developed for the experimental part of this paper is publicly available at https://github.com/gliuzzi/GMM.
Acknowledgments
The authors are very grateful to Prof. Luigi Grippo for the useful discussions at the early stages of this work.

### v3

@@
 To the best of our knowledge, these results are novel to the literature.
-Moreover, extensive computational experiments show that the gradient methods with momentum here presented outperform classical conjugate gradient methods and are (at least) competitive with the state-of-art method for unconstrained optimization, i.e, L-BFGS method.
+Moreover, extensive computational experiments show that the gradient method with momentum here presented is a solid choice to tackle some classes of nonconvex unconstrained problems.
 keywords:
@@
 Furthermore, we provide complexity results, proving for the proposed gradient method with momentum the worst-case complexity bound of ${\cal O}(\epsilon^{-2})$, which is optimal for first order algorithms in the nonconvex setting [18].
-Up to our knowledge, the presented algorithm is the first framework of gradient methods with momentum having in the nonconvex case both theoretical convergence properties and a complexity bound.
-Extensive computational experiments show that the gradient methods with momentum here presented outperform conjugate gradient methods and are (at least) competitive with the state-of-art method for unconstrained optimization, i.e, L-BFGS method.
+Up to our knowledge, the theoretical results presented here are novel in the literature.
+Known convergence and complexity properties for similar algorithmic frameworks are indeed based on assumptions that are stronger than ours or, as an alternative, can only be enforced in practice resorting to restarting techniques.
+Our analysis here does not make hypotheses concerning second order information for the objective nor linear independence between gradient and momentum directions; in addition, the conditions resulting from our analysis can be rigorously handled directly in the subspace, which is a crucial aspect from the computational perspective.
+Extensive numerical experiments show that the gradient methods with momentum here presented outperform popular implementations of the conjugate gradient method and are (at least) competitive with a widespread limited-memory Quasi-Newton method like L-BFGS algorithm [19].
+Moreover, while the conjugate gradient based procedure implemented in cg_descent software [20, 21] proves to still remain superior, the proposed method seems comparable, with some upsides, in the particular, relevant class of multi-layer perceptron training problems [22, 23].
 The rest of the paper is organized as follows: we describe the main idea of the work in section 2, with a focus on some important related works in section 2.1; in section 3 we discuss conditions for the proposed method to be well defined.
@@
 $\min_{d\in\mathbb{R}^{n}}\;\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}B_{k}d,$ (2)
-where $B_{k}$ is a suitable symmetric positive definite matrix. By properly choosing $B_{k}$, we retrieve standard methods; in particular:
+where $B_{k}$ is a suitable symmetric positive definite matrix. By properly choosing $B_{k}$, we retrieve standard methods, in particular:
   • if $B_{k}=I$, then we obtain the steepest descent direction $-\nabla f(x_{k})$;
@@
 The quadratic Taylor polynomial for $\psi_{k}(\alpha,\beta)$ centered at $(0,0)$ is thus
-$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (10)
+$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (10)
 The matrix $H_{k}$ defined by (6) can then be viewed as a matrix approximating $\nabla^{2}\psi_{k}(0,0)$.
 From this point of view, $H_{k}$ could in fact be any $2\times 2$ matrix (independent of the $n\times n$ matrix $B_{k}$) related to the approximation of the quadratic Taylor polynomial of $\psi_{k}(\alpha,\beta)=f(x_{k}-\alpha g_{k}+\beta s_{k})$, i.e.,
-$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (11)
+$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (11)
 Setting $u=[\alpha\;\;\beta]^{T}$, the minimization problem of (11) leads again to problem (8), being $H_{k}$ a generic $2\times 2$ matrix with suitable properties.
-From a theoretical point of view, the following issues must me considered:
+From a theoretical point of view, the following issues must be considered:
   1. (a)
@@
   2. (b)
-    we shall analyze under which conditions the obtained search direction $d_{k}$, coupled with line search techniques, allows to prove global convergence properties and complexity results for the iterative scheme (9).
+    we shall analyze under which conditions the obtained search direction $d_{k}$, coupled with line search techniques, allow to prove global convergence properties and complexity results for the iterative scheme (9).
 In the following, we will state conditions concerning either the $n\times n$ matrix $B_{k}$ appearing in (7) or directly the $2\times 2$ matrix $H_{k}$ in (8) (with no explicit connection to $B_{k}$), in order to satisfy the above theoretical requirements.
-2.1 Related Works
-We now briefly analyze some important works directly related to our approach.
-Subspace minimization CG
-Subspace minimization conjugate gradient (SMCG) methods, for which we refer the reader, for example, to the very recent paper [8], consider the two-dimensional subproblem (4)-(5) by assuming that the $n\times n$ matrix $B_{k}$ satisfies the secant equation
+2.1 Related works and open problems
+Before turning to the theoretical analysis, we shall briefly discuss some important works directly related to the approach addressed in this paper.
+Subspace minimization conjugate gradient (SMCG) methods
+The class of SMCG type algorithms, for which we refer the reader, for example, to the very recent paper [8], consider the two-dimensional subproblem (4)-(5) by assuming that the $n\times n$ matrix $B_{k}$ satisfies the secant equation
 $B_{k}s_{k}=g_{k}-g_{k-1}.$
@@
 However, as clearly written in [8], “A question is naturally to be asked: can one develop an efficient SMCG method without determining the parameter $\rho_{k}$?”.
+In other words, is it possible to define a convergent SMCG method without the need of handling the full matrix $B_{k}$?
 By our work, we give a positive answer to this question.
@@
 Convergence results are stated by assuming that $\{B_{k}\}$ is a sequence of uniformly positive definite symmetric matrices bounded above.
-Then, gradient methods with momentum fits this framework provided that $g_{k}$ and $s_{k}$ are linearly independent.
-The convergence analysis relies on the properties of the $n\times n$ matrix $B_{k}$.
+However, while the convergence analysis relies on the properties of the $n\times n$ matrix $B_{k}$, it is crucial, for computational performance, to only deal with the small matrix $H_{k}$.
 (Riemannian) dimension-reduced second-order method (RDRSOM)
@@
 This assumption can not be relaxed since the proposed method is based on the cubic regularization approach.
-Our main theoretical contribution, with respect to all the works on SMCG methods and to [19], concerns the definition of conditions on the $2\times 2$ matrix $H_{k}$ sufficient to ensure that the sequence of search directions $\{d_{k}\}$ is gradient-related and hence to state convergence and complexity results for the proposed gradient methods with momentum.
+Our main theoretical contribution with respect to all the works on SMCG methods and to [24], thus concerns the definition of conditions on the $2\times 2$ matrix $H_{k}$ sufficient to ensure that the sequence of search directions $\{d_{k}\}$ is gradient-related and hence to state convergence and complexity results for the proposed gradient methods with momentum.
+Note that working with only the $2\times 2$ matrix, regardless of the $n\times n$ matrix $B_{k}$, is fundamental to make these algorithmic frameworks computationally viable at large scale; our result on this issue thus enlightens how it can be done in a theoretically sound manner.
 As regards the contribution with respect to [26], we do not state assumptions on the Hessian matrix of the objective function: we only require that the gradient is Lipschitz continuous.
 Thus, we attain the same complexity result under much weaker assumptions.
-We remark that the theoretical focus on the low-dimensional matrix $H_{k}$ is fundamental to design convergent algorithms for large-scale optimization regardless of the $n\times n$ matrix $B_{k}$.
-This issue yields relevant implications, deeply analyzed and discussed later, from theoretical, algorithmic and computational point of views.
 3 Existence of a solution for the two dimensional problem
@@
 $x_{k+1}=x_{k}+\eta_{k}d_{k},$
-where $\eta_{k}$ is the stepsize computed by an Armjio-type line search, is globally convergent to stationary points, Moreover, under this assumption the algorithm can be proved to have a worst case iteration and function evaluations complexity bound of ${\cal O}(\epsilon^{-2})$ to reach a solution $\bar{x}$ with $\|\nabla f(\bar{x})\|\leq\epsilon$.
+where $\eta_{k}$ is the stepsize computed by an Armjio-type line search, is globally convergent to stationary points.
+Moreover, under this assumption the algorithm can be proved to have a worst case iteration and function evaluations complexity bound of ${\cal O}(\epsilon^{-2})$ to reach a solution $\bar{x}$ with $\|\nabla f(\bar{x})\|\leq\epsilon$.
 Note that this bound is actually tight for first-order methods under standard first-order smoothness assumptions.
@@
 holds for all $k$. Then the direction $d_{k}$ satisfies the following conditions:
-$g_{k}^{T}d_{k}\leq-\frac{\eta_{1}}{\eta_{2}^{2}}\|g_{k}\|^{2}$ (15)
-$\frac{1}{\eta_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{\eta_{1}}\|g_{k}\|.$ (16)
+$g_{k}^{T}d_{k}\leq-\frac{\eta_{1}}{\eta_{2}^{2}}\|g_{k}\|^{2}$
+$\frac{1}{\eta_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{\eta_{1}}\|g_{k}\|.$
 Proof.
@@
 By the next proposition, we finally state suitable conditions on the sequence of matrices $\{H_{k}\}$, that also take into account the sequences of vectors involved, i.e., $\{g_{k}\}$ and $\{s_{k}\}$.
-These conditions are sufficient to ensure that the obtained sequence of search directions $\{d_{k}\}$ is gradient-related.
+These conditions represent the main theoretical contribution of the paper, as they are sufficient to ensure that the obtained sequence of search directions $\{d_{k}\}$ is gradient-related.
 Proposition 5.
@@
 $g_{k}^{T}d_{k}\leq-\frac{1}{c_{2}}\|g_{k}\|^{2}$ (28)
-$\frac{1}{c_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{c_{1}}\|g_{k}\|.$ (30)
+$\frac{1}{c_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{c_{1}}\|g_{k}\|.$
 Proof.
@@
 $H_{k}=\left[\begin{array}[]{cc}\|g_{k}\|(\hat{H}_{11})_{k}\|g_{k}\|&\|g_{k}\|(\hat{H}_{12})_{k}\|s_{k}\|\\ \|s_{k}\|(\hat{H}_{21})_{k}\|g_{k}\|&\|s_{k}\|(\hat{H}_{22})_{k}\|s_{k}\|\end{array}\right].$
+We will show later in this work how the condition can practically be enforced in computational scenarios.
 5 Algorithmic model
-In this section we propose an algorithmic framework for gradient methods with momentum which exploits the theoretical analysis carried out in the previous sections.
+In this section we present an algorithmic framework for gradient methods with momentum which exploits the theoretical analysis carried out in the previous sections.
 The idea is to define a first-order algorithm with both strong theoretical guarantees in the nonconvex setting and the possibility of computationally exploiting eventual second-order information on the minimization problem.
@@
 From a computational perspective, this allows to possibly save several backtracking steps and, consequently, function evaluations.
-1 Input: $x_{0}\in\mathbb{R}^{n}$, $\gamma\in(0,1)$, $\delta\in(0,1)$, $c_{1}>0$, $c_{2}>0$.
-2 Set $k\leftarrow 0$
-3 while $\|\nabla f(x_{k})\|>0$ do
-       /* Compute the search direction */
-4       Set gr_dir_found $\leftarrow$ False
-5       Define a $2\times 2$ symmetric matrix $H_{k}$
-6       if problem (37) admits solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ then
-7             Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
-8              if $g_{k}^{T}d_{k}\leq-c_{1}\|g_{k}\|^{2}$ and $\|d_{k}\|\leq c_{2}\|g_{k}\|$  then
-9                    Set gr_dir_found $\leftarrow$ True
-10
-11      if gr_dir_found $=$ False then
-12             Define a new $2\times 2$ symmetric matrix $H_{k}$ satisfying condition (28)
-13              Compute $\alpha_{k}$ and $\beta_{k}$ by solving (37)
-14             Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
-15
-      /* Perform Armijo line search along $d_{k}$/
-16
-17      Set $\eta\leftarrow 1$
-18       while $f(x_{k}+\eta d_{k})>f(x_{k})+\gamma\eta d_{k}^{\top}\nabla f(x_{k})$ do
-19              Set $\eta\leftarrow\delta\eta$
-20
-21      Set $\eta_{k}\leftarrow\eta$, $x_{k+1}\leftarrow x_{k}+\eta_{k}d_{k}$
-22       Set $k\leftarrow k+1$
 Algorithm 1 Gradient Method with Momentum (GMM)
+1:Input: $x_{0}\in\mathbb{R}^{n}$, $\gamma\in(0,1)$, $\delta\in(0,1)$, $c_{1}>0$, $c_{2}>0$.
+2:Set $k\leftarrow 0$
+3:while $\|\nabla f(x_{k})\|>0$ do
+4:     /$\ast$/ ∗ Compute the search direction $\ast$/∗ /
+5:     Set gr_dir_found $\leftarrow$ False
+6:     Define a $2\times 2$ symmetric matrix $H_{k}$
+7:     if problem (35) admits solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ then
+8:         Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
+9:         if $g_{k}^{T}d_{k}\leq-c_{1}\|g_{k}\|^{2}$ and $\|d_{k}\|\leq c_{2}\|g_{k}\|$  then
+10:              Set gr_dir_found $\leftarrow$ True
+11:     if gr_dir_found $=$ False then
+12:         Define a new $2\times 2$ symmetric matrix $H_{k}$ satisfying condition (26)
+13:         Compute $\alpha_{k}$ and $\beta_{k}$ by solving (35)
+14:         Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
+15:     /$\ast$/ ∗ Perform Armijo line search along $d_{k}$ $\ast$/∗ /
+16:     Set $\eta\leftarrow 1$
+17:     while $f(x_{k}+\eta d_{k})>f(x_{k})+\gamma\eta d_{k}^{\top}\nabla f(x_{k})$ do
+18:         Set $\eta\leftarrow\delta\eta$
+19:     Set $\eta_{k}\leftarrow\eta$, $x_{k+1}\leftarrow x_{k}+\eta_{k}d_{k}$
+20:     Set $k\leftarrow k+1$
 The theoretical properties of the proposed framework for gradient methods with momentum are stated in the following proposition.
@@
 From the steps the algorithm and Proposition 5 we have that $\{d_{k}\}$ is a sequence of gradient-related directions.
-Since the Armijo line search is employed within Algorithm 1, the results follow from [24, Proposition 1.2.1] and by [23].
+Since the Armijo line search is employed within Algorithm 1, the results follow from [29, Proposition 1.2.1] and by [28].
 ∎
-We now focus on the issue of modifying a given $2\times 2$ matrix $H_{k}$ in order to satisfy condition (28) as required at step 1 of Algorithm 1.
-Suppose that a symmetric $H_{k}$ matrix has been defined at step 1, but either problem problem (37) does not admit solution or the test at step 1 is not satisfied, i.e., the value of gr_dir_found remains False.
-Then, step 1 must be performed and a new matrix $H_{k}$ must be constructed modifying as least as possible the given matrix defined at step 1.
-Let us denote by $H_{k}^{0}$ the matrix defined at step 1 and by $H_{k}$ the new matrix defined at step 1. We can proceed as follows:
-    Let $\hat{H}_{k}$ be obtained by a modified Cholesky factorization [24] applied to ${D_{k}^{-1}}H_{k}^{0}D_{k}^{-1}$, where
-    $D_{k}=\begin{bmatrix}\|g_{k}\|&0\\ 0&\|s_{k}\|\end{bmatrix};$
-    Set $H_{k}=D_{k}\hat{H}_{k}D_{k}$, i.e.,
+We now focus on the issue of modifying a given $2\times 2$ matrix $H_{k}$ in order to satisfy condition (26) as required at step 12 of Algorithm 1.
+Suppose that a symmetric $H_{k}$ matrix has been defined at step 6, but either problem (35) does not admit solution or the test at step 9 is not satisfied, i.e., the value of gr_dir_found remains False.
+Then, step 12 must be performed and a new matrix $H_{k}$ must be constructed modifying as least as possible the given matrix defined at step 6.
+Let us denote by $H_{k}^{0}$ the matrix defined at step 6 and by $H_{k}$ the new matrix defined at step 12. We can proceed as follows:
+    A matrix $\hat{H}_{k}$ can be obtained by a modified Cholesky factorization [29] applied to $H_{k}^{0}$
+    Letting
+    $D_{k}=\begin{bmatrix}\|g_{k}\|&0\\ 0&\|s_{k}\|\end{bmatrix}$
+    we can set $H_{k}=D_{k}\hat{H}_{k}D_{k}$, i.e.,
     $H_{k}=\left[\begin{array}[]{cc}\|g_{k}\|(\hat{H}_{11})_{k}\|g_{k}\|&\|g_{k}\|(\hat{H}_{12})_{k}\|s_{k}\|\\ \|s_{k}\|(\hat{H}_{21})_{k}\|g_{k}\|&\|s_{k}\|(\hat{H}_{22})_{k}\|s_{k}\|\end{array}\right].$
 The boundedness of $\{H_{k}^{0}\}$ and the properties of the modified Cholesky factorization imply that (34) of Remark 2 holds, so that, according to the same remark, the matrix $H_{k}$ is such that condition (26) is satisfied.
-We conclude this section by stating a theoretical result showing that, under suitable assumptions on the objective function and a choice of the matrix $H_{k}$ actually related to the Hessian $\nabla^{2}f(x_{k})$, steps 1-1 are never executed for $k$ sufficiently large, that is, the test at step 1 is always satisfied.
+We conclude this section by stating a theoretical result showing that, under suitable assumptions on the objective function and a choice of the matrix $H_{k}$ actually related to the Hessian $\nabla^{2}f(x_{k})$, steps 12-14 are never executed for $k$ sufficiently large, that is, the test at step 9 is always satisfied.
 Proposition 7.
@@
 $H_{k}=\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}$
-is the matrix defined at step 1, and let $\{x_{k}\}$ be the sequence generated by Algorithm 1.
+is the matrix defined at step 6, and let $\{x_{k}\}$ be the sequence generated by Algorithm 1.
 Assume that $\{x_{k}\}$ converges to $x^{*}$, where $\nabla f(x^{*})=0$ and the Hessian matrix $\nabla^{2}f(x^{*})$ is positive definite.
-Furthermore, assume that the constants of the test at step 1 are chosen in a such a way that:
+Furthermore, assume that the constants of the test at step 9 are chosen in a such a way that:
 $c_{1}\leq\theta^{3}\frac{\lambda_{\min}(\nabla^{2}f(x^{*}))}{\lambda_{\max}(\nabla^{2}f(x^{*}))^{2}},\qquad\qquad c_{2}\geq\frac{2}{\theta\lambda_{\min}(\nabla^{2}f(x^{*}))}.$
-where $\theta\in(0,1)$. Then, for $k$ sufficiently large the test at step 1 is satisfied.
+where $\theta\in(0,1)$. Then, for $k$ sufficiently large the test at step 9 is satisfied.
 Proof.
@@
 6.1 Approximating Hessian-vector products by finite differences of gradients
-In line with the strategy employed in [20, 21], we can draw inspiration by the Truncated Newton methods approach [25, 26], where the explicit management of the Hessian matrix $\nabla^{2}f(x_{k})$ is not required, but rather the Hessian-vector product $\nabla^{2}f(x_{k})d$ is directly handled, with $d\in R^{n}$.
+In line with the strategy employed in [25, 26], we can draw inspiration by the strategy employed in Truncated Newton methods [30, 31], where the explicit management of the Hessian matrix $\nabla^{2}f(x_{k})$ is not required, but rather the Hessian-vector product $\nabla^{2}f(x_{k})d$ is directly handled, with $d\in R^{n}$.
 Consider $H_{k}$ defined by (6), i.e.,
-$H_{k}=\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix},$
+$H_{k}=\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix}.$
 The elements of $H_{k}$ can be obtained estimating the two matrix-vector products $B_{k}g_{k}$ and $B_{k}s_{k}$ by finite difference approximation, namely by the following vectors (where $\xi>0$ is a suitably small parameter):
@@
 Assuming that $f$ is twice continuously differentiable, we have seen — see (10) — that the quadratic Taylor polynomial of $\psi_{k}(\alpha,\beta)$ centered at $(0,0)$ is
-$f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$
+$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$
 Then, the matrix $H_{k}$ defined by (6) can be viewed as a matrix approximating $\nabla^{2}\psi_{k}(0,0)$, and, in a more general setting, $H_{k}$ could be any $2\times 2$ matrix (independent of the $n\times n$ matrix $B_{k}$).
@@
 The described strategy requires three function evaluations. However, considering the point $(\alpha_{1},\beta_{1})=(0,-1)$, we have
-$f(x_{k}-\alpha_{1}g_{k}+\beta_{1}s_{k})=f(x_{k}-(x_{k}-x_{k-1})=f(x_{k-1})$
+$f(x_{k}-\alpha_{1}g_{k}+\beta_{1}s_{k})=f(x_{k}-(x_{k}-x_{k-1})=f(x_{k-1}).$
 Then, by exploiting the information of the past iteration, i.e., the knowledge of $f(x_{k-1})$, the matrix $H_{k}$ can be built by only two additional function evaluations.
@@
 In this section, we describe and report the results of thorough computational experiments aimed at assessing the potential of the algorithm proposed in this work.
-Specifically, we considered to this aim a benchmark of 163 problems from the CUTEst test-suite [28].
+We carried out the tests on two different sets of problems; the results of the two cases are presented in two different subsections.
+The code for all the experiments described in this work was written in Python 3.11, exploiting numpy, scipy and pytorch libraries. The software is available at: https://github.com/gliuzzi/GMM.
+Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rules discussed in section 6; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq.
+26 holds.
+As for the line search, we set $\delta=0.5,\;\gamma=10^{-5}$.
+The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration.
+As a baseline for comparisons, we took into account classical nonlinear conjugate gradient methods [33] and the L-BFGS algorithm [19], the latter arguably being the most popular solver for the solution of smooth unconstrained nonlinear optimization problems.
+For both algorithms, we considered the very popular implementations available through the scipy library.
+In a subsequent stage of our experimentation, we also took into account the cg_descent procedure available at https://people.clas.ufl.edu/hager/software/ and based on [34, 20, 33, 21].
+The cg_descent provides a highly efficient implementation of a conjugate-gradient type method with subspace optimization steps.
+We used the python wrapper of the original C code of cg_descent v.
+6.8, available at https://github.com/alexfikl/pycgdescent, so that we could employ the procedure within the same testing environment as the other considered methods.
+The results of the experiments are shown in the manuscript in the form of performance profiles [35].
+The detailed tables with all the result can be accessed in the code repository https://github.com/gliuzzi/GMM.
+7.1 CUTEst problems
+We considered a benchmark of 168 problems from the CUTEst test-suite [36].
 In particular, we included in our test set all 183 problems in the CUTEst collection that are 1) unconstrained, 2) regular, 3) with a number of variables $n$ that is greater or equal than 1000 (except for the problems LUKSAN12LS, LUKSAN13LS, LUKSAN14LS having 998 variables) or that is user definable; in the latter case, we set the number of variables to 1000 or as close as possible to 1000.
@@
 Those problems are marked in table 1 with a strike-through.
-The code for all the experiments described in this work was written in Python 3.9, only exploiting numpy and scipy libraries. The software is available at:
-Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rules discussed in section 6; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq.
-28 holds.
-As for the parameters, we set $\delta=0.5,\;\gamma=10^{-5}$.
-The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration.
-As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-3}$; we also set a maximum number of iterations to 5000, so that we consider a failure each run that ends by hitting this threshold.
-As a baseline for comparisons, we took into account classical nonlinear conjugate gradient methods [29] and the L-BFGS algorithm [30], which can arguably be considered the state-of-the-art for the solution of smooth unconstrained nonlinear optimization problems.
-For both algorithms, we considered the very efficient implementations available through the scipy library.
-The results of the experiments are shown in the form of performance profiles [31]. As performance metrics, we considered both the runtime and the number of iterations.
-In the following, we report results for three versions of our algorithm, depending on the technique adopted for the computation of $H_{k}$, as detailed in section 6.
+As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-6}$; we also set a maximum number of iterations to 5000, so that we consider a failure each run that ends by hitting this threshold.
+As performance metrics, we considered both the runtime and the number of iterations.
+In the following, we first report results for three versions of our algorithm, depending on the technique adopted for the computation of $H_{k}$, as detailed in section 6.
 More in detail, we denote by GMM1, GMM2, GMM3 the algorithm where $H_{k}$ is computed as described in subsection 6.1, 6.2 and 6.3, respectively.
-Figure 1: Performance profiles on all 163 problems.
+We also included in this first test scipy conjugate gradient implementation, denoted by CGscipy, as a preliminary baseline.
+Figure 1: Performance profiles on all 168 problems.
 Figure 2: Performance profiles on the subset of problems where all the solvers find the same best function value.
 In fig.
-1, we report the performance profiles in terms of time and number of iterations for the comparison among GMM1, GMM2, GMM3 and CGscipy on the whole collection of 163 problems.
-As we can see, the best performing solver is GMM2 both in terms of time and iterations.
-We can also say that GMM3 is the second best solver in terms of time even though it is the worst one in terms of iterations.
-Both solvers GMM2 and GMM3 are better than CGscipy in terms of time.
-In doing these considerations, we must also take into account that it might happen that solvers converge toward different points.
-Indeed, in table 2 we report the number of wins, i.e., the number of problems where each solver finds the best function value.
-Table 2: Number of wins, i.e., the number of problems where each solver finds the best function value, on all 163 test problems. GMM variants and CGscipy solvers are considered.
-Solver # wins
-GMM1 71
-GMM2 58
-GMM3 26
-CGscipy 16
+1, we report the performance profiles in terms of time and number of iterations for this comparison comparison among GMM1, GMM2, GMM3 and CGscipy on the whole collection of 168 problems.
+As we can see, the best performing solver is GMM2 both in terms of time and, together with GMM1 also in terms of iterations.
+Both solvers GMM2 and GMM3 are also better than CGscipy in terms of time.
 Then, in fig. 2, we report the performance profiles considering the subset of problems where each solver finds the same best function value, i.e.
-$f^{*}_{i}-f_{L}<10^{-3}\quad\forall\ i\in\cal S,$
+$f^{*}_{i}-f_{L}<10^{-6}\quad\forall\ i\in\cal S,$
 where $\cal S$ is the set of solvers.
 From fig.
-2, we note that the performance of GMM1 becomes closer to that of the conjugate gradient method in terms of computational time, whereas GMM2 and GMM3 continue to be clearly superior.
-Figure 3: Performance profiles for the comparison between GMM1, GMM2, GMM3 and L-BFGS on all the 163 problems.
-Figure 4: Performance profiles for the comparison between GMM1, GMM2, GMM3 and L-BFGS on the subset of problems where all the solvers find the same best function value.
-Now, we consider the comparison between our methods and the state-of-the-art algorithm for unconstrained optimization, namely, L-BFGS [30].
-In particular, performance profiles of solvers GMM1, GMM2, GMM3 and L-BFGS on the entire collection of 163 problems are reported in fig.
+2, we note that the performance of GMM methods continues to be clearly superior to that of the baseline conjugate gradient method in terms of computational time, especially in the case of GMM2.
+Based on the above results, we will be considering version GMM2 in the rest of the Section as the choice for our GMM algorithm.
+In what follows, we also allowed a slight degree of non-monotoniticy in GMM, based on the rule proposed in [37].
+We then turned to the comparison between our method and some state-of-the-art solvers for unconstrained optimization, namely, L-BFGS [19] and the cg_descent procedure [34, 20, 21].
+Performance profiles of solvers GMM and L-BFGS on the entire collection of 168 problems are reported in fig.
 3.
-As we can see, our method GMM2 is quite competitive with L-BFGS in terms of computational time.
-Again, we have to consider the fact that the different solvers can converge toward points with different function values.
-In fact, as reported in table 3, GMM1 and GMM2 find a better function value more frequently than L-BFGS.
-Table 3: Number of wins, i.e., the number of problems where each solver finds the best function value, on all 163 test problems. $GMM$ variants and L-BFGS solvers are considered.
-Solver # wins
-GMM1 55
-GMM2 40
-GMM3 17
-L-BFGSscipy 33
-In fig.
-4, we thus consider the performance profiles obtained for the four solvers on the subset of problems where all of them find the same function value.
-As we can see, GMM2 gained some efficiency and it now slightly dominates L-BFGS in terms of computational time.
-Then, for major clarity, we report in figs. 5 and 6 the performance profiles for the comparison between our bet method GMM2 and L-BFGS.
-Figure 5: Performance profiles for the comparison between GMM2 and L-BFGS on all the 163 problems.
-Figure 6: Performance profiles for the comparison between GMM2 and L-BFGS on the subset of problems where all the solvers find the same best function value.
-From figs.
-5 and 6, we can say that GMM2 is quite competitive if not better than L-BFGS both in terms of computational time and number of iterations.
-The advantages of using GMM2 are more evident when the two solvers are compared on the subset of problems where they both find the same function value.
+As we can see, GMM seems to be slightly superior to L-BFGS in both metrics.
+Figure 3: Performance profiles for the comparison between GMM and L-BFGS on all the 168 problems.
+The picture heavily changes if cg_descent is included in the comparative analysis: as we can see from fig.
+5, it appears by far superior than the other solvers considered in the experimentation on this benchmark of problems.
+From the one hand, the superior depth and quality of implementation of the cg_descent procedure w.r.t.
+the simple python code implementing the GMM is probably a component that makes the gap so particularly pronounced; on the other hand, although clearly outperformed, GMM still retains a certain degree of comparability with the state-of-the-art method.
+This conclusion made us interested in carrying out a further comparison on a different and relevant class of problems.
+Figure 4: Performance profiles for the comparison between GMM, cg-descent and L-BFGS on all the 168 problems.
+7.2 Multi-layer perceptron training problems
+As an additional test for the proposed GMM method we considered a set of problems from supervised learning context, where methods with momentum are particularly popular.
+In particular, we aim to train multi-layer perceptrons for binary classification tasks [22, 23].
+We considered a fully connected networks consisting of two hidden layers of 128 units each.
+The input layer matches the size of the samples of each dataset.
+The output layer consists in a single unit with sigmoid activation.
+The activation function of the hidden units is the hyperbolic tangent.
+In total, 30 binary classification tasks are considered.
+We limit the dimension of the training set by taking at most 10000 samples from each dataset.
+We do so by randomly drawing the samples from the full dataset, as this allows to maintain the class imbalance of some datasets.
+A recap of the problems considered is available in table 2.
+Table 2: Details of the 30 binary classification problems considered. Value $n$ denotes the number of parameters for the problem.
+Dataset $n$ Dataset $n$
+a1a 32001 gisette 656769
+a2a 32001 ijcnn1 19585
+a3a 32385 liver.disorders 17409
+a4a 32385 phishing 25473
+a5a 32385 sonar 24449
+a6a 32385 svmguide1 17281
+a7a 32513 svmguide3 19585
+a8a 32513 w1a 55169
+a9a 32513 w2a 55169
+breast.cancer 18049 w3a 55169
+cod.rna 17793 w4a 55169
+covtype 23681 w5a 55169
+diabetes 17793 w6a 55169
+fourclass 17025 w7a 55169
+german.numer 19841 w8a 55169
+The weights in the multi-layer perceptron are initialized using Kaiming uniform initialization [38].
+We then ran both GMM and cg_descent from the same starting point, repeating this process with three different random seeds.
+The objective function is binary cross-entropy.
+We also added an $\ell_{2}$ regularization term, with a weight $\lambda_{\textbf{reg}}=0.001$.
+The termination condition is $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-3}$, which is a typical accuracy degree for learning tasks.
+We report in fig.
+5 the performance profiles in terms of time, number of iterations and value of the objective function obtained for GMM and cg_descent.
+We considered each pair dataset-seed as a distinct problem in the profiles, so that we have 90 instances overall.
+Figure 5: Performance profiles for the comparison between GMM and cg-descent on the multi-layer perceptron training problems.
+We can see from these plots that cg_descent is once again superior in terms of runtime; however, we do not observe this time differences in the number of instances solved; moreover, in a quarter of the instances GMM finished faster, and the overall performance gap appears more moderate.
+Actually, if we look at the number of iterations the two solver really seem comparable - the difference in runtime is arguably due to the higher number of function evaluations required by our GMM.
+Another element of interest in these results concerns the final objective function value attained by the two solvers; GMM seems to have a slight advantage in this regard; we find this aspect intriguing, as as the momentum term is often used in machine learning also to escape bad stationary points in the highly nonconvex landscapes resulting from training problems.
 8 Conclusions
 In this work, we introduced a general framework of gradient methods with momentum for nonconvex optimization.
-For the proposed class of algorithms, we proved global convergence and optimal worst-case complexity bounds.
+For the proposed class of algorithms, we proved global convergence and optimal worst-case complexity bound.
 The assumptions required to obtain the theoretical results are lighter to check, from a computational perspective, than those required in related works from the literature.
 This result allowed to devise particularly efficient ways of implementing the proposed method.
-Thorough computational results we showed that the novel algorithm outperforms standard conjugate gradient methods and is competitive with L-BFGS, the state-of-the-art for nonlinear optimization problems.
-Future studies might concern improvements of the proposed framework and suitable modifications to make it exploitable in different settings.
+By thorough computational results we showed that a rather simple implementation of the novel algorithm outperforms simpler versions of conjugate gradient methods, is competitive with L-BFGS, and is somewhat comparable to state-of-the-art solvers for nonlinear optimization problems, especially in particular classes of problems, such as multi-layer perceptron training tasks.
+These positive preliminary feedback on our procedure motivates future studies aimed at the further improvement of the performance of our procedure.
+In particular, the integration of some limited-memory and subspace optimization subroutines in cg_descent within the GMM framework might be subject of future studies.
 Declarations
@@
 Acknowledgments
-The authors are very grateful to Prof. Luigi Grippo for the useful discussions at the early stages of this work.
+The authors are very grateful to Prof.
+Luigi Grippo for the useful discussions at the early stages of this work.
+We are also extremely thankful to the Editor-in-Chief and the anonymous referees whose very stimulating comments allowed us to improve the soundness of our work.

### v4

@@
 To the best of our knowledge, these results are novel to the literature.
-Moreover, extensive computational experiments show that the gradient method with momentum here presented is a solid choice to tackle some classes of nonconvex unconstrained problems.
+Moreover, extensive computational experiments show that the gradient method with momentum here presented is competitive with respect to other popular solvers for nonconvex unconstrained problems.
 keywords:
@@
 Our analysis here does not make hypotheses concerning second order information for the objective nor linear independence between gradient and momentum directions; in addition, the conditions resulting from our analysis can be rigorously handled directly in the subspace, which is a crucial aspect from the computational perspective.
-Extensive numerical experiments show that the gradient methods with momentum here presented outperform popular implementations of the conjugate gradient method and are (at least) competitive with a widespread limited-memory Quasi-Newton method like L-BFGS algorithm [19].
-Moreover, while the conjugate gradient based procedure implemented in cg_descent software [20, 21] proves to still remain superior, the proposed method seems comparable, with some upsides, in the particular, relevant class of multi-layer perceptron training problems [22, 23].
+Extensive numerical experiments show that the gradient methods with momentum here presented outperform popular implementations of the conjugate gradient method and a widespread limited-memory Quasi-Newton method like L-BFGS algorithm [19].
+Moreover, while the state-of-the-art conjugate gradient based procedure implemented in cg_descent software [20, 21] proves to be clearly superior, the proposed method may represent an interesting baseline for devising a solver competitive with the state of the art.
 The rest of the paper is organized as follows: we describe the main idea of the work in section 2, with a focus on some important related works in section 2.1; in section 3 we discuss conditions for the proposed method to be well defined.
@@
 where $B_{k}$ is a suitable symmetric positive definite matrix. By properly choosing $B_{k}$, we retrieve standard methods, in particular:
-  • if $B_{k}=I$, then we obtain the steepest descent direction $-\nabla f(x_{k})$;
-  • if $B_{k}=\nabla^{2}f(x_{k})$, then we obtain Newton’s direction $-\left[\nabla^{2}f(x_{k})\right]^{-1}\nabla f(x_{k})$;
+  • if $B_{k}=I$, then we obtain the steepest descent direction $\nabla f(x_{k})$;
+  • if $B_{k}=\nabla^{2}f(x_{k})$, then we obtain Newton’s direction $\left[\nabla^{2}f(x_{k})\right]^{-1}\nabla f(x_{k})$;
   • if $B_{k}$ is a positive definite matrix obtained with suitable update rules, we obtain standard Quasi-Newton updates.
@@
 or equivalently
-$\phi(\alpha,\beta)=\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (5)
+$\phi(\alpha,\beta)=\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}B_{k}g_{k}&-g_{k}^{T}B_{k}s_{k}\\ -g_{k}^{T}B_{k}s_{k}&s_{k}^{T}B_{k}s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}$.
+(5)
 We can denote the $2\times 2$ matrix in the above equation as
@@
 The quadratic Taylor polynomial for $\psi_{k}(\alpha,\beta)$ centered at $(0,0)$ is thus
-$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (10)
+$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}g_{k}^{T}\nabla^{2}f(x_{k})g_{k}&-g_{k}^{T}\nabla^{2}f(x_{k})s_{k}\\ -g_{k}^{T}\nabla^{2}f(x_{k})s_{k}&s_{k}^{T}\nabla^{2}f(x_{k})s_{k}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}$.
+(10)
 The matrix $H_{k}$ defined by (6) can then be viewed as a matrix approximating $\nabla^{2}\psi_{k}(0,0)$.
 From this point of view, $H_{k}$ could in fact be any $2\times 2$ matrix (independent of the $n\times n$ matrix $B_{k}$) related to the approximation of the quadratic Taylor polynomial of $\psi_{k}(\alpha,\beta)=f(x_{k}-\alpha g_{k}+\beta s_{k})$, i.e.,
-$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (11)
+$\psi_{k}(\alpha,\beta)=f(x_{k})+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}\begin{bmatrix}H_{11}&H_{12}\\ H_{12}&H_{22}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}$.
+(11)
 Setting $u=[\alpha\;\;\beta]^{T}$, the minimization problem of (11) leads again to problem (8), being $H_{k}$ a generic $2\times 2$ matrix with suitable properties.
@@
 Proposition 3.
-Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of (7).
+Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of (7).
 Let $\{B_{k}\}\subseteq\mathbb{R}^{n\times n}$ be the sequence of symmetric matrices defining problems (7) and assume that there exist scalars $0<\eta_{1}\leq\eta_{2}$ such that
@@
 being $d_{k}=P_{k}u_{k}$. Now, consider the first row of system (15), i.e.,
-$-g_{k}^{T}B_{k}d_{k}=\|g_{k}\|^{2}.$
+$g_{k}^{T}B_{k}d_{k}=\|g_{k}\|^{2}.$
 Using (16) and recalling assumption (14), we can write
@@
 Recalling (16), using (14) and (17), it follows
-$-g_{k}^{T}d_{k}\geq\lambda_{\min}(B_{k})\|d_{k}\|^{2}\geq\eta_{1}\|d_{k}\|^{2}\geq{\frac{\eta_{1}}{\eta_{2}^{2}}}\|g_{k}\|^{2}.$ (18)
+$g_{k}^{T}d_{k}\geq\lambda_{\min}(B_{k})\|d_{k}\|^{2}\geq\eta_{1}\|d_{k}\|^{2}\geq{\frac{\eta_{1}}{\eta_{2}^{2}}}\|g_{k}\|^{2}.$ (18)
 Considering again (16) we can also write
@@
 Then, proposition 3 shows that condition eq.
-14 on the eigenvalues of the matrices $B_{k}$ guarantees that the solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ of (8) produces directions $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$ that are gradient-related.
+14 on the eigenvalues of the matrices $B_{k}$ guarantees that the solution $\alpha_{k}\;\;\beta_{k}]^{T}$ of (8) produces directions $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$ that are gradient-related.
 However, checking or ensuring that a sequence of $n\times n$ matrices $\{B_{k}\}$ have uniformly bounded eigenvalues can be extremely challenging for large-dimensional optimization problems.
 Problem (8) can be rewritten in the following form:
-$\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}H_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (20)
+$\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}H_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}$.
+(20)
 In this section we analyze conditions on the $2\times 2$ matrix $H_{k}$ sufficient to guarantee suitable properties to the direction $d_{k}$.
@@
 Proposition 4.
-Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k},$ where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is the solution of (20).
+Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k},$ where $\alpha_{k}\;\;\beta_{k}]^{T}$ is the solution of (20).
 Let $\{H_{k}\}\subseteq\mathbb{R}^{2\times 2}$ be the sequence of symmetric matrices defining problems (20) and assume that there exist scalars $0<\hat{c}_{1}\leq\hat{c}_{2}$ such that
@@
 $g_{k}^{T}d_{k}\leq-\frac{1}{\hat{c}_{2}}\|g_{k}\|^{4}$ (23)
-$\frac{1}{\hat{c}_{2}}\|g_{k}\|^{3}\leq\|d_{k}\|\leq\frac{1}{\hat{c}_{1}}(\|g_{k}\|+\|s_{k}\|)(\|g_{k}\|^{2}+|g_{k}^{T}s_{k}|).$ (23)
+$\frac{1}{\hat{c}_{2}}\|g_{k}\|^{3}\leq\|d_{k}\|\leq\frac{1}{\hat{c}_{1}}(\|g_{k}\|+\|s_{k}\|)(\|g_{k}\|^{2}+|g_{k}^{T}s_{k}|).$
 Proof.
@@
 $\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}=\begin{bmatrix}\|g_{k}\|^{2}&-g_{k}^{T}s_{k}\end{bmatrix}H_{k}^{-1}\begin{bmatrix}\|g_{k}\|^{2}\\ -g_{k}^{T}s_{k}\end{bmatrix}$. (25)
-We note that $-g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$; thus, by using eq. 21 and (25) we have:
-$-g_{k}^{T}d_{k}\geq\lambda_{min}(H^{-1}_{k})\biggl{(}\|g_{k}\|^{4}+(g_{k}^{T}s_{k})^{2}\biggr{)},$
+We note that $g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$; thus, by using eq. 21 and (25) we have:
+$g_{k}^{T}d_{k}\geq\lambda_{min}(H^{-1}_{k})\biggl(\|g_{k}\|^{4}+(g_{k}^{T}s_{k})^{2}\biggr),$
 which implies:
@@
 $D_{k}=\begin{bmatrix}\|g_{k}\|&0\\ 0&\|s_{k}\|\end{bmatrix}$.
-Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $[\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of problem (20). Then, the direction $d_{k}$ satisfies the following conditions:
+Let $d_{k}$ be defined by $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$, where $\alpha_{k}\;\;\beta_{k}]^{T}$ is solution of problem (20). Then, the direction $d_{k}$ satisfies the following conditions:
 $g_{k}^{T}d_{k}\leq-\frac{1}{c_{2}}\|g_{k}\|^{2}$ (27)
-$\frac{1}{c_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{c_{1}}\|g_{k}\|.$
+$\frac{1}{c_{2}}\|g_{k}\|\leq\|d_{k}\|\leq\frac{2}{c_{1}}\|g_{k}\|.$ (28)
 Proof.
@@
 Multiplying both sides of the above equation by $\begin{bmatrix}\|g_{k}\|&-\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}$ we obtain:
-$\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}=\begin{bmatrix}\|g_{k}\|&-\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}\tilde{H}_{k}^{-1}\begin{bmatrix}\|g_{k}\|\\ -\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}.$ (30)
-Recalling eq. 26 and that $-g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$, we have from (30):
-$-g_{k}^{T}d_{k}\geq\lambda_{min}(\tilde{H}^{-1})\biggl{(}\|g_{k}\|^{2}+\left(\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\right)^{2}\biggr{)},$
+$\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}=\begin{bmatrix}\|g_{k}\|&-\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}\tilde{H}_{k}^{-1}\begin{bmatrix}\|g_{k}\|\\ -\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\end{bmatrix}$.
+(30)
+Recalling eq. 26 and that $g_{k}^{T}d_{k}=\alpha_{k}\|g_{k}\|^{2}-\beta_{k}g_{k}^{T}s_{k}$, we have from (30):
+$g_{k}^{T}d_{k}\geq\lambda_{min}(\tilde{H}^{-1})\biggl(\|g_{k}\|^{2}+\left(\frac{g_{k}^{T}s_{k}}{\|s_{k}\|}\right)^{2}\biggr),$
 and thus
@@
     define a quadratic subproblem
-    $\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}H_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}.$ (35)
+    $\min_{\alpha,\beta}$ $\frac{1}{2}\ \begin{bmatrix}\alpha\\ \beta\end{bmatrix}^{T}H_{k}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}+\begin{bmatrix}-\|g_{k}\|^{2}\\ g_{k}^{T}s_{k}\end{bmatrix}^{T}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}$.
+(35)
     where $H_{k}$ is a $2\times 2$ symmetric matrix;
   • (b)
-    once computed a solution $[\alpha_{k}\ \beta_{k}]^{T}$ of (35), provided one exists, a check on the obtained search direction
+    once computed a solution $\alpha_{k}\ \beta_{k}]^{T}$ of (35), provided one exists, a check on the obtained search direction
     $d_{k}=-\alpha_{k}g_{k}+\beta_{k}s_{k}$
@@
 3:while $\|\nabla f(x_{k})\|>0$ do
-4:     /$\ast$/ ∗ Compute the search direction $\ast$/∗ /
+[…]
 5:  Set gr_dir_found $\leftarrow$ False
 6:  Define a $2\times 2$ symmetric matrix $H_{k}$
-7:     if problem (35) admits solution $[\alpha_{k}\;\;\beta_{k}]^{T}$ then
+7:  if problem (35) admits solution $\alpha_{k}\;\;\beta_{k}]^{T}$ then
 8:   Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
@@
 10:     Set gr_dir_found $\leftarrow$ True
-11:     if gr_dir_found $=$ False then
+11:  if gr_dir_found == False then
 12:   Define a new $2\times 2$ symmetric matrix $H_{k}$ satisfying condition (26)
@@
 14:   Set $d_{k}\leftarrow-\alpha_{k}g_{k}+\beta_{k}s_{k}$
-15:     /$\ast$/ ∗ Perform Armijo line search along $d_{k}$ $\ast$/∗ /
+[…]
 16:  Set $\eta\leftarrow 1$
@@
 From the steps the algorithm and Proposition 5 we have that $\{d_{k}\}$ is a sequence of gradient-related directions.
-Since the Armijo line search is employed within Algorithm 1, the results follow from [29, Proposition 1.2.1] and by [28].
+Since the Armijo line search is employed within Algorithm 1, the results follow from [27, Proposition 1.2.1] and by [26].
 ∎
@@
 In this section, we describe and report the results of thorough computational experiments aimed at assessing the potential of the algorithm proposed in this work.
-We carried out the tests on two different sets of problems; the results of the two cases are presented in two different subsections.
-The code for all the experiments described in this work was written in Python 3.11, exploiting numpy, scipy and pytorch libraries. The software is available at: https://github.com/gliuzzi/GMM.
-Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rules discussed in section 6; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq.
+The code for all the experiments described in this work was written in Python 3.11, exploiting numpy and scipy libraries. The software is available at: https://github.com/gliuzzi/GMM.
+Regarding Algorithm 1, it has been implemented choosing matrix $H_{k}$ according to the rule discussed in section 6.2, which provided better results than those from section 6.1 and section 6.3 in preliminary experiments; we employed the safeguarding technique based on the modified Cholesky factorization, described in section 5, to ensure that Assumption eq.
 26 holds.
 As for the line search, we set $\delta=0.5,\;\gamma=10^{-5}$.
+We also allowed a slight degree of non-monotoniticy in the line search, based on the rule proposed in [31].
 The point $x_{-1}$ is set equal to $x_{0}$, so that the momentum term is null at the first iteration.
@@
 For both algorithms, we considered the very popular implementations available through the scipy library.
-In a subsequent stage of our experimentation, we also took into account the cg_descent procedure available at https://people.clas.ufl.edu/hager/software/ and based on [34, 20, 33, 21].
+We also took into account the state-of-the-art cg_descent procedure available at https://people.clas.ufl.edu/hager/software/ and based on [33, 20, 32, 21].
 The cg_descent provides a highly efficient implementation of a conjugate-gradient type method with subspace optimization steps.
@@
 The detailed tables with all the result can be accessed in the code repository https://github.com/gliuzzi/GMM.
-7.1 CUTEst problems
-We considered a benchmark of 168 problems from the CUTEst test-suite [36].
-In particular, we included in our test set all 183 problems in the CUTEst collection that are 1) unconstrained, 2) regular, 3) with a number of variables $n$ that is greater or equal than 1000 (except for the problems LUKSAN12LS, LUKSAN13LS, LUKSAN14LS having 998 variables) or that is user definable; in the latter case, we set the number of variables to 1000 or as close as possible to 1000.
-In Table 1 those 183 problems are reported along with their dimensions.
-Table 1: Collection of 183 unconstrained CUTEst problems ($n$ denotes the number of variables).
-The problems whose name is crossed out have then been removed from the initial selection as they turned out to be unbounded from below.
+We considered a benchmark of 77 unconstrained problems from the CUTEst test-suite [35] with a number of variables $n\geq 1000$.
+In Table 1 these 77 problems are listed along with their dimension; note that, for problems with a user definable number of variables we always choose the default value.
+Table 1: Collection of 77 unconstrained CUTEst problems ($n$ denotes the number of variables).
 Problem $n$ Problem $n$ Problem $n$
-ARGLINA 1000 EDENSCH 2000 OBSTCLBU 9604
-ARGLINB 1000 EIGENALS 2550 ODC 4900
-ARGLINC 1000 EIGENBLS 2550 ODNAMUR 11130
-ARGTRIGLS 1000 EIGENCLS 2652 OSCIGRAD 100000
-BDEXP 5000 ENGVAL1 5000 OSCIPATH 5000
-BOX 10000 EXTROSNB 1000 PENALTY1 1000
-BOXPOWER 10000 FLETBV3M 5000 PENALTY2 1000
-BRATU1D 5001 FLETCBV2 5000 PENALTY3 1000
-BROWNAL 1000 FLETCBV3 5000 PENTDI 5000
-BROYDN3DLS 5000 FLETCHBV 5000 POWELLBC 1000
-BROYDNBDLS 5000 FLETCHCR 1000 POWELLSG 5000
-CHEBYQAD 1000 FMINSRF2 5625 POWER 10000
-CHNRSNBM 1000 FMINSURF 5625 POWERSUM 1000
-CLPLATEA 4970 FREUROTH 5000 PRICE3 2000
-CURLY30 10000 GENHUMPS 5000 PROBPENL 1000
-CVXBQP1 10000 GENROSE 1000 QING 1000
-CYCLIC3LS 100002 GENROSEB 1000 QRTQUAD 5000
-CYCLOOCFLS 29996 GRIDGENA 5560 QUARTC 5000
-CYCLOOCTLS 29996 HILBERTA 1000 RAYBENDL 2046
-DEGDIAG 100001 HILBERTB 1000 RAYBENDS 2046
-DEGTRID 100001 INDEF 5000 S368 1000
-ARWHEAD 5000 INDEFM 100000 SBRYBND 5000
-BDQRTIC 5000 INTEQNELS 1002 SCHMVETT 5000
-BROYDN7D 5000 JNLBRNG1 9604 SCOSINE 5000
-BRYBND 5000 JNLBRNG2 9604 SCURLY10 10000
-CHAINWOO 4000 JNLBRNGA 9604 SCURLY20 10000
-CLPLATEB 4970 JNLBRNGB 9604 SCURLY30 10000
-CLPLATEC 4970 KSSLS 1000 SENSORS 1000
-COSINE 10000 LIARWHD 5000 SINEALI 1000
-CRAGGLVY 5000 LINVERSE 1999 SINQUAD 5000
-CURLY10 10000 LMINSURF 5329 SPARSINE 5000
-CURLY20 10000 LUKSAN11LS 1000 SPARSQUR 10000
-DEGTRID2 100001 LUKSAN12LS 998 SPIN2LS 1002
-DIAGIQB 1000 LUKSAN13LS 998 SPINLS 1327
-DIAGIQE 1000 LUKSAN14LS 998 SROSENBR 5000
-DIAGIQT 1000 LUKSAN15LS 1000 SSBRYBND 5000
-DIAGNQB 1000 LUKSAN16LS 1000 SSC 4900
-DIAGNQE 1000 LUKSAN17LS 1000 SSCOSINE 5000
-DIAGNQT 1000 LUKSAN21LS 1000 STRTCHDV 1000
-DIAGPQB 1000 LUKSAN22LS 1000 TESTQUAD 5000
-DIAGPQE 1000 MANCINO 1000 TOINTGSS 5000
-DIAGPQT 1000 MCCORMCK 5000 TORSION1 5184
-DIXMAANA1 3000 MODBEALE 20000 TORSION2 5184
-DIXMAANB 3000 MOREBV 5000 TORSION3 5184
-DIXMAANC 3000 NCB20 5010 TORSION4 5184
-DIXMAAND 3000 NCB20B 5000 TORSION5 5184
-DIXMAANE1 3000 NCVXBQP1 10000 TORSION6 5184
-DIXMAANF 3000 NCVXBQP2 10000 TORSIONA 5184
-DIXMAANG 3000 NCVXBQP3 10000 TORSIONB 5184
-DIXMAANH 3000 NLMSURF 5329 TORSIONC 5184
-DIXMAANI1 3000 NOBNDTOR 5184 TORSIOND 5184
-DIXMAANJ 3000 NONCVXU2 5000 TORSIONE 5184
-DIXMAANK 3000 NONCVXUN 5000 TORSIONF 5184
-DIXMAANL 3000 NONDIA 5000 TQUARTIC 5000
-DIXMAANM1 3000 NONDQUAR 5000 TRIDIA 5000
-DIXMAANN 3000 NONMSQRT 4900 TRIGON1 1000
-DIXMAANO 3000 NONSCOMP 5000 TRIGON2 1000
-DIXMAANP 3000 OBSTCLAE 9604 VANDANMSLS 1002
-DIXON3DQ 10000 OBSTCLAL 9604 VARDIM 1000
-DQDRTIC 5000 OBSTCLBL 9604 VAREIGVL 5000
-DQRTIC 5000 OBSTCLBM 9604 WOODS 4000
-Then, from this set of 183 problems, we removed those problems that are unbounded below, i.e., the problems where at least one solver among the ones considered generates a sequence $\{x_{k}\}$ such that $f(x_{k})\to-\infty$.
-Those problems are marked in table 1 with a strike-through.
-As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-6}$; we also set a maximum number of iterations to 5000, so that we consider a failure each run that ends by hitting this threshold.
+ARWHEAD 5000 DIXMAANM1 3000 MSQRTBLS 1024
+BDQRTIC 5000 DIXMAANN 3000 NCB20 5010
+BOX 10000 DIXMAANO 3000 NCB20B 5000
+BOXPOWER 20000 DIXMAANP 3000 NONCVXU2 5000
+BROYDN3DLS 5000 DIXON3DQ 10000 NONDIA 5000
+BROYDN7D 5000 DQDRTIC 5000 NONDQUAR 5000
+BROYDNBDLS 5000 DQRTIC 5000 PENALTY1 1000
+BRYBND 5000 EDENSCH 2000 POWELLSG 5000
+CHAINWOO 4000 EG2 1000 POWER 10000
+COSINE 10000 EIGENALS 2550 QUARTC 5000
+CRAGGLVY 5000 EIGENBLS 2550 SCHMVETT 5000
+CURLY10 10000 EIGENCLS 2652 SINQUAD 5000
+CURLY20 10000 ENGVAL1 5000 SINQUAD2 5000
+CURLY30 10000 EXTROSNB 1000 SPARSINE 5000
+DIXMAANA1 3000 FLETBV3M 5000 SPARSQUR 10000
+DIXMAANB 3000 FLETCBV2 5000 SPINLS 1327
+DIXMAANC 3000 FLETCHCR 1000 SPMSRTLS 4999
+DIXMAAND 3000 FMINSRF2 5625 SROSENBR 5000
+DIXMAANE1 3000 FMINSURF 5625 SSBRYBND 5000
+DIXMAANF 3000 FREUROTH 5000 SSCOSINE 5000
+DIXMAANG 3000 GENHUMPS 5000 TOINTGSS 5000
+DIXMAANH 3000 JIMACK 3549 TQUARTIC 5000
+DIXMAANI1 3000 KSSLS 1000 TRIDIA 5000
+DIXMAANJ 3000 LIARWHD 5000 VAREIGVL 5000
+DIXMAANK 3000 MOREBV 5000 WOODS 4000
+DIXMAANL 3000 MSQRTALS 1024
+As stopping condition for our algorithm, as well as all other methods considered in the experimentation, we required $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-6}$; we also set a maximum number of iterations to 1000000, so that we consider a failure each run that ends by hitting this threshold.
 As performance metrics, we considered both the runtime and the number of iterations.
-In the following, we first report results for three versions of our algorithm, depending on the technique adopted for the computation of $H_{k}$, as detailed in section 6.
-More in detail, we denote by GMM1, GMM2, GMM3 the algorithm where $H_{k}$ is computed as described in subsection 6.1, 6.2 and 6.3, respectively.
-We also included in this first test scipy conjugate gradient implementation, denoted by CGscipy, as a preliminary baseline.
-Figure 1: Performance profiles on all 168 problems.
-Figure 2: Performance profiles on the subset of problems where all the solvers find the same best function value.
-In fig.
-1, we report the performance profiles in terms of time and number of iterations for this comparison comparison among GMM1, GMM2, GMM3 and CGscipy on the whole collection of 168 problems.
-As we can see, the best performing solver is GMM2 both in terms of time and, together with GMM1 also in terms of iterations.
-Both solvers GMM2 and GMM3 are also better than CGscipy in terms of time.
-Then, in fig. 2, we report the performance profiles considering the subset of problems where each solver finds the same best function value, i.e.
-$f^{*}_{i}-f_{L}<10^{-6}\quad\forall\ i\in\cal S,$
-where $\cal S$ is the set of solvers.
-From fig.
-2, we note that the performance of GMM methods continues to be clearly superior to that of the baseline conjugate gradient method in terms of computational time, especially in the case of GMM2.
-Based on the above results, we will be considering version GMM2 in the rest of the Section as the choice for our GMM algorithm.
-In what follows, we also allowed a slight degree of non-monotoniticy in GMM, based on the rule proposed in [37].
-We then turned to the comparison between our method and some state-of-the-art solvers for unconstrained optimization, namely, L-BFGS [19] and the cg_descent procedure [34, 20, 21].
-Performance profiles of solvers GMM and L-BFGS on the entire collection of 168 problems are reported in fig.
-3.
-As we can see, GMM seems to be slightly superior to L-BFGS in both metrics.
-Figure 3: Performance profiles for the comparison between GMM and L-BFGS on all the 168 problems.
-The picture heavily changes if cg_descent is included in the comparative analysis: as we can see from fig.
-5, it appears by far superior than the other solvers considered in the experimentation on this benchmark of problems.
-From the one hand, the superior depth and quality of implementation of the cg_descent procedure w.r.t.
+The results of our experiments are shown in Figure 1.
+As we can see, cg_descent procedure appears by far superior than all the other solvers on the considered benchmark of problems.
+On the other hand, GMM outperforms both L-BFGS and scipy’s CG method.
+The superior depth and quality of implementation of the cg_descent procedure w.r.t.
 the simple python code implementing the GMM is probably a component that makes the gap so particularly pronounced; on the other hand, although clearly outperformed, GMM still retains a certain degree of comparability with the state-of-the-art method.
-This conclusion made us interested in carrying out a further comparison on a different and relevant class of problems.
-Figure 4: Performance profiles for the comparison between GMM, cg-descent and L-BFGS on all the 168 problems.
-7.2 Multi-layer perceptron training problems
-As an additional test for the proposed GMM method we considered a set of problems from supervised learning context, where methods with momentum are particularly popular.
-In particular, we aim to train multi-layer perceptrons for binary classification tasks [22, 23].
-We considered a fully connected networks consisting of two hidden layers of 128 units each.
-The input layer matches the size of the samples of each dataset.
-The output layer consists in a single unit with sigmoid activation.
-The activation function of the hidden units is the hyperbolic tangent.
-In total, 30 binary classification tasks are considered.
-We limit the dimension of the training set by taking at most 10000 samples from each dataset.
-We do so by randomly drawing the samples from the full dataset, as this allows to maintain the class imbalance of some datasets.
-A recap of the problems considered is available in table 2.
-Table 2: Details of the 30 binary classification problems considered. Value $n$ denotes the number of parameters for the problem.
-Dataset $n$ Dataset $n$
-a1a 32001 gisette 656769
-a2a 32001 ijcnn1 19585
-a3a 32385 liver.disorders 17409
-a4a 32385 phishing 25473
-a5a 32385 sonar 24449
-a6a 32385 svmguide1 17281
-a7a 32513 svmguide3 19585
-a8a 32513 w1a 55169
-a9a 32513 w2a 55169
-breast.cancer 18049 w3a 55169
-cod.rna 17793 w4a 55169
-covtype 23681 w5a 55169
-diabetes 17793 w6a 55169
-fourclass 17025 w7a 55169
-german.numer 19841 w8a 55169
-The weights in the multi-layer perceptron are initialized using Kaiming uniform initialization [38].
-We then ran both GMM and cg_descent from the same starting point, repeating this process with three different random seeds.
-The objective function is binary cross-entropy.
-We also added an $\ell_{2}$ regularization term, with a weight $\lambda_{\textbf{reg}}=0.001$.
-The termination condition is $\|\nabla f(x_{k})\|_{\infty}\leq 10^{-3}$, which is a typical accuracy degree for learning tasks.
-We report in fig.
-5 the performance profiles in terms of time, number of iterations and value of the objective function obtained for GMM and cg_descent.
-We considered each pair dataset-seed as a distinct problem in the profiles, so that we have 90 instances overall.
-Figure 5: Performance profiles for the comparison between GMM and cg-descent on the multi-layer perceptron training problems.
-We can see from these plots that cg_descent is once again superior in terms of runtime; however, we do not observe this time differences in the number of instances solved; moreover, in a quarter of the instances GMM finished faster, and the overall performance gap appears more moderate.
-Actually, if we look at the number of iterations the two solver really seem comparable - the difference in runtime is arguably due to the higher number of function evaluations required by our GMM.
-Another element of interest in these results concerns the final objective function value attained by the two solvers; GMM seems to have a slight advantage in this regard; we find this aspect intriguing, as as the momentum term is often used in machine learning also to escape bad stationary points in the highly nonconvex landscapes resulting from training problems.
+This consideration makes us believe that the integration of preconditioning or subspace minimization steps in GMM might lead to an easily accessible Python code with a comparable efficiency w.r.t.
+cg_descent.
+Figure 1: Performance profiles for the comparison between GMM, cg_descent and the Scipy’s implementations of the conjugate gradient method and L-BFGS, on the 77 problems considered.
 8 Conclusions